### v2

# A Review of Hydrogen-Enabled Resilience Enhancement for Multi-Energy Systems

Abstract
Ensuring resilience in multi-energy systems (MESs) becomes both more urgent and more challenging due to the rising occurrence and severity of extreme events (e.g., natural disasters, extreme weather, and cyber-physical attacks). Among many measures of strengthening MES resilience, the integration of hydrogen shows exceptional potential in cross-temporal flexibility, cross-spatial flexibility, cross-sector flexibility, and black start capability. Although many hydrogen-enabled MES resilience enhancement measures have been developed, the current literature lacks a systematic overview of hydrogen-enabled resilience enhancement in MESs. To fill the research gap, this paper provides a comprehensive overview of hydrogen-enabled MES resilience enhancement. First, advantages and challenges of adopting hydrogen in MES resilience enhancement are summarized. Then, we propose a resilience enhancement framework for hydrogen-enabled MESs. Under the proposed framework, existing resilience metrics and event-oriented contingency models are summarized and discussed. Furthermore, we classify hydrogen-enabled planning measures by the types of hydrogen-related facilities and provide some insights for planning problem formulation frameworks. Moreover, we categorize the hydrogen-enabled operation enhancement measures into three operation response stages: preventive, emergency, and restoration. Finally, we identify some research gaps and point out possible future directions in aspects of comprehensive resilience metric design, temporally-correlated event-targeted scenario generation, multi-type temporal-spatial cyber-physical contingency modeling under compound extreme events, multi-network multi-timescale coordinated planning and operation, low-carbon resilient planning and operation, and large language model-assisted whole-process resilience enhancement.
Index Terms:
Multi-energy systems, hydrogen-enabled resilience enhancement, resilience metrics, contingency modeling, planning, operation, extreme events, low-carbon
I Introduction
It is well recognized that power and energy systems are indispensable to modern society[1]. Compared with single-energy systems, multi-energy systems (MESs), which integrate electricity, heat, gas, and other energy carriers, can offer more advantages in terms of efficiency, flexibility, economic performance, decarbonization, and energy security. Thus, MESs play an important role in shaping our energy future[2]. However, the operation of MESs is facing growing challenges due to the rising frequency and intensity of extreme events, such as natural disasters (e.g., earthquakes, floods, wildfires), climate-induced extreme weather (e.g., heatwaves, hurricanes, cold spells), and cyber-physical attacks targeting critical infrastructure[3, 4, 5, 6]. These events can significantly affect interconnected energy supply networks, and their correlated occurrence across time and space (e.g., concurrent heatwaves and cyber threats) often leads to compound risks with amplified socio-economic impacts[7]. For instance, blackout-related economic losses from extreme weather events are estimated at 20–55 billion dollars per year in the United States[4]. To address these challenges, the concept of resilience has been introduced, emphasizing an energy system’s ability to withstand, adapt to, and recover from extreme events while maintaining essential functions[5], [8, 9, 10]. As MESs couple multiple energy carriers and play a critical role in supporting both critical loads and economic activities, ensuring their resilience is not only a technical necessity but also a societal imperative.
Among the measures of enhancing MES resilience, the integration of hydrogen energy resources (HERs) has shown exceptional potential. To be specific, HERs refer broadly to the various technologies and systems that produce, store, transport, and utilize hydrogen as an energy carrier, including stationary HERs (e.g., electrolyzers, hydrogen tanks, hydrogen pipelines, hydrogen fuel cell-based conversion systems, and hydrogen refueling stations (HRSs)) and mobile HERs (MHERs), which include hydrogen tube trailers (HTTs)[11, 12, 13, 14, 15, 16, 17], mobile hydrogen trucks[18, 19, 20], mobile fuel cell trucks[21], hydrogen fuel cell buses (HFCBs)[22], hydrogen fuel cell vehicles (HFCVs)[23]. The above HERs can offer distinguishing features for enhancing MES resilience in aspects of energy supply and carbon emission[24, 25, 26, 27], i.e., cross-temporal flexibility, cross-spatial flexibility, cross-sector flexibility, and black start capability. Despite the promising potential of hydrogen integration, developing effective resilience enhancement strategies for hydrogen-enabled MESs (HMESs) involves significant complexity across multiple dimensions. First, defining resilience metrics for HMESs is inherently challenging due to the need to capture many aspects, e.g., load loss, carbon resilience, and physical and functional features of hydrogen systems (e.g., hydrogen availability). Second, modeling extreme event-related HESS contingencies requires careful representation of multi-stage processes (such as electrolysis, compression, storage, and reconversion) and each process is subject to distinct cyber and physical operational risks. Third, designing the schemes of hydrogen-related infrastructure planning and HMES operation faces solving challenges caused by uncertainties, mixed integer variables, multi-timescale variables, temporal-spatial coupling constraints, nonlinear and nonconvex constraints, multi-energy coupling constraints, and multi-objective function.
To overcome the above challenges, researchers have defined many resilience metrics, developed some contingency models, and proposed many planning and operation methods for enhancing the resilience of HMESs[28, 29, 30, 31, 32]. However, there is a lack of a systematic overview of hydrogen-enabled resilience enhancement for MESs. To fill the research gap, we provide a comprehensive overview on leveraging hydrogen to enhance the resilience for MESs. Although there have been many review works related to power grid resilience enhancement[34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 33] with different perspectives (e.g., networked microgrids, event timeline of enhancement, cyber-physical interdependence, complex network, and cross-domain multilayer architecture), they neglect to consider the HERs. [33] is the first and only work to review the resilience of a hydrogen-powered smart grid. Although this review paper considers the hydrogen-related facilities, there are many differences between [33] and our work in aspects of objects, perspectives, contents, and research gaps. Firstly, we focus on HMESs, while [33] focuses on hydrogen-enabled smart grid. Secondly, we review the existing works from the perspective of hydrogen-enabled planning and operation. In contrast, [33] mainly focuses on the event timeline of measures, i.e., pre-event, during-event, and post-event. Thirdly, this paper summarizes advantages and challenges of adopting hydrogen in MES resilience enhancement, resilience metrics of HMES and event-oriented contingency models, which are not mentioned in [33]. Fourthly, we classify the existing resilience enhancement measures with the consideration of hydrogen pipeline hardening, more types of HERs, and multi-stage operation coordination, whereas [33] considers measures only within independent stages. Finally, we identify six research gaps and point out future directions, which are neglected in [33]. To the best of our knowledge, our work is the first survey dedicated to hydrogen-enabled resilience enhancement for MESs.
The main contributions of this paper can be summarized below.
  1. 1.
    We summarize the advantages and challenges of adopting hydrogen in MES resilience enhancement. Specifically, distinguishing features of hydrogen for resilience enhancement, resilience potential and security issues of hydrogen-related technologies, and optimization challenges of hydrogen-enabled planning and operation enhancement are summarized.
  2. 2.
    We propose a comprehensive resilience enhancement framework for HMESs considering hydrogen-enabled planning and operation measures (including pre-event preventive response, during-event emergency response, and post-event restoration response). Under the proposed framework, the widely used resilience metrics and event-oriented contingency models in existing works are summarized.
  3. 3.
    We classify the planning enhancement of HMES resilience from two aspects, i.e., single-type hydrogen facility and multi-type hydrogen facility. The considered hydrogen-related facilities consist of HESSs, hydrogen pipelines, HRSs, HFCBs, fuel cell-driven combined cooling, heat, and power (CCHP), and hydrogen-related production facilities. Moreover, the insights in aspects of uncertainty representation and formulation framework for HMES planning are summarized.
  4. 4.
    We categorize the operation enhancement of HMES resilience according to three types of operation response stages involved, including preventive response, emergency response, and restoration response. Moreover, the insights about the particular function of hydrogen in each operation stage are summarized.
  5. 5.
    We identify six research gaps and point out future directions in aspects of comprehensive resilience metric design, temporally-correlated event-targeted scenario generation, multi-type temporal-spatial cyber-physical contingency modeling under compound extreme events, multi-network multi-timescale coordinated planning and operation, low-carbon resilient planning and operation, and large language model-assisted whole-process resilience enhancement.
The rest of this paper is organized as follows. In Section II, advantages and challenges of adopting hydrogen in MES resilience enhancement are summarized. In Section III, a comprehensive resilience enhancement framework for HMESs is described. In Section IV, the definitions and metrics of HMES resilience are described. In Section V, event-oriented contingency modeling is summarized. In Section VI, planning enhancement measures for HMES resilience are classified, and some insights are provided. In Section VII, operation enhancement measures for HMES resilience are categorized. Section VIII identifies six research gaps and possible future directions. Finally, conclusions and lessons learned are summarized in Section IX. For better understanding, we provide the list of abbreviations in alphabetical order in Table I.
TABLE I: List of abbreviations in alphabetical order
Abbreviation Description
C&CG Column and constraint generation
CCHP Combined cooling, heat and power
CRI Comprehensive resilience index
ELNS Expected load not served
ERL Expected restored load
ESS Energy storage system
GSHP Ground-source heat pump
H2P Hydrogen-to-power
HESS Hydrogen energy storage system
HER Hydrogen energy resource
HFCB Hydrogen fuel cell bus
HFCV Hydrogen fuel cell vehicle
HMES Hydrogen-enabled multi-energy system
HRS Hydrogen refueling station
HTT Hydrogen tube trailer
LLM Large language model
$L$ Load loss
LSR Load served ratio
MHER Mobile hydrogen energy resource
MILP Mixed integer linear programming
MINLP Mixed integer nonlinear programming
MISOCP Mixed-integer second-order cone programming
PV Photovoltaic
SHS Seasonal hydrogen storage
TES Thermal energy storage
WSLL Weighted sum of load loss
II Advantages and Challenges of Adopting Hydrogen in MES Resilience Enhancement
In this section, we summarize the advantages and challenges of adopting hydrogen in MES resilience enhancement. To be specific, the distinguishing features of hydrogen for MES resilience enhancement are firstly provided. Then, from the perspective hydrogen supply chain, resilience potential and security issues related to hydrogen-related technologies are summarized. Finally, the optimization advantages in planning enhancement and operation enhancement of MES resilience are analyzed.
Figure 1: Distinguishing features of hydrogen in MES resilience enhancement
II-A Distinguishing Features of Hydrogen for MES Resilience Enhancement
Among various measures of enhancing MES resilience, the integration of HERs has shown exceptional potential due to their distinctive multi-functional attributes as shown in Fig. 1, which are described below.
II-A1 Cross-temporal Flexibility
Hydrogen can provide cross-temporal flexibility[46, 47, 48]. Unlike traditional battery energy storage systems, hydrogen can be stored at a much larger scale and over longer durations, e.g., underground salt-cavern hydrogen storage can be used to provide cross-seasonal low-cost and large-scale hydrogen storage[49]. The above feature can improve MES energy resilience by providing energy backup in the case of long-duration power outages. For example, electricity and heat/cold demand in the summer and winter are typically higher than those in spring and autumn. To prepare for possible prolonged outages (e.g., lasting for more than 48 hours[50, 51, 52]) under extreme events (e.g., heat waves or ice disaster), proactive hydrogen production and storage could be conducted in spring and autumn. In addition, when produced from renewable energy, hydrogen improves the consumption rate of renewable energy. By converting excess electricity into storable hydrogen using seasonal energy storages (SHSs) and conventional HESSs, MESs can buffer supply-demand mismatches more effectively and operate with greater carbon resilience (i.e., an energy system’s ability to maintain or recover to its current and predicted emission level during and after shocks such as renewable intermittency, public health emergencies, or climate-related disruptions)[53]. As shown in [54], 100% self-sufficiency and zero carbon emission can be simultaneously achieved in distant oceanic islands with abundant renewable energy resources by coordinating SHSs and HESSs.
II-A2 Cross-spatial Flexibility
Hydrogen can provide cross-spatial flexibility[46, 47, 48]. For example, hydrogen can be flexibly transported across regions through hydrogen pipelines, HTTs, or liquid hydrogen trucks, making it possible to allocate energy geographically according to regional demand and supply conditions. This spatial transfer capability becomes especially valuable in resilience enhancement when local renewable resources are insufficient or disrupted. For instance, during an earthquake, hydrogen can be transported from unaffected industrial production bases to disaster-hit areas to power temporary shelters via fuel cells, ensuring heating and lighting during prolonged grid outages. Similarly, mobile hydrogen refueling trucks can be deployed to support remote or damaged areas after extreme weather events, such as floods, by providing temporary hydrogen fuel for microgrids and emergency vehicles. By bridging spatial gaps between supply and demand, hydrogen transportation infrastructure allows MESs to dynamically reallocate resources, maintain critical services, and prevent cascading failures in interconnected energy networks, resulting in an improved energy resilience.
II-A3 Cross-sector Flexibility
Hydrogen plays a crucial role in enhancing the interoperability of energy systems by acting as a versatile bridge between multiple energy carriers, including electricity, heat, gas, and transport. For example, the power-to-hydrogen technologies enable surplus electricity from renewable sources like wind and solar to be converted into hydrogen, which can then be stored and later used for electricity and heat generation or as a fuel for HFCVs in transportation sector. In addition, hydrogen can be converted to synthetic methane or blended with natural gas for transportation[31],[55],[56], which can reduce dependence on pure natural gas. The above interoperability not only enhances system efficiency, but also supports more robust and adaptive energy supply chains. Based on dynamic reconfiguration, hydrogen can facilitate resource sharing across different energy vectors, especially during extreme events. In such cases, if one subsystem experiences a failure, hydrogen can provide an alternative energy pathway, thereby ensuring the continuous operation of essential services and improving the MES resilience.
II-A4 Black Start Capability
Hydrogen plays a critical role in safeguarding key infrastructures (e.g., hospitals, emergency shelters, and communication networks) during extreme events. In the event of prolonged power outages or natural gas supply failures, hydrogen-based infrastructure (such as proton exchange membrane fuel cells and hydrogen tanks) can operate off-grid and provide autonomous “black start” capabilities[57], significantly enhancing the system’s ability to self-recover without external intervention during extreme events. According to [57], black start should be implemented before network reconfiguration and load restoration could be done. Compared with black-start diesel generators, hydrogen fuel cells produce zero local emissions, operate quietly, and offer higher reliability with fewer maintenance requirements.
In summary, hydrogen has some distinguishing features and its role in enhancing energy resilience and carbon resilience of MESs is multi-dimensional, which can be summarized as shown in Table II.
TABLE II: Distinguishing features of hydrogen and its role in MES resilience enhancement
Index Distinguishing Feature Role in MES Resilience Enhancement Energy Resilience Carbon Resilience
1 Cross-temporal flexibility Ability to store hydrogen over long durations (e.g., seasonal storage) and shift energy use from surplus periods to shortage periods Long-duration backup and seasonal balancing reduce risk of supply shortages Enables higher renewable integration and reduces curtailment
2 Cross-spatial flexibility Capability to transport hydrogen across regions via pipelines, HTTs, or liquid hydrogen trucks to match spatially uneven supply-demand patterns Allows resource reallocation during local outages or regional shortages Enables renewable-rich regions to export clean energy to high-demand or low-renewable areas
3 Cross-sector flexibility Interoperability between electricity, heat, gas, and transport sectors via hydrogen conversion and blending Provides alternative energy pathways when one sector fails Facilitates green fuel substitution and decarbonization in multiple sectors
4 Black start capability Off-grid operation of hydrogen fuel cells and tanks to restart or sustain critical infrastructure during outages Ensures rapid recovery and maintains essential services in blackout scenarios limited enhancement
II-B Resilience Potential of Hydrogen-related Technologies
In the hydrogen supply chain, there are many hydrogen production pathways (e.g., electrolyzer, or steam methane reforming [58]), hydrogen storage manners (e.g., salt-cavern hydrogen storage, hydrogen tank), transportation methods (e.g., pipelines, trucks, or tube trailers), and hydrogen-to-power technologies (e.g., proton exchange membrane fuel cells, solid oxide fuel cells, and turbines), which offer different resilience capabilities in aspects of black start capability, off-grid operability, and response time, as summarized in Table III. By coordinating different hydrogen-related technologies, synergies can be achieved. For example, by combining proton exchange membrane fuel cell and on-site hydrogen tanks, fast, clean, and autonomous black start could be implemented for key infrastructures (e.g., hospitals). electrolyzers, salt cavern storage, pipelines, and hydrogen turbines, large-scale grid restoration could be implemented. Thus, a careful selection of hydrogen-related technologies tailored to the specific application scenarios is required for designing hydrogen-enabled MESs.
TABLE III: Resilience Potential of Hydrogen-related Technologies
Category Technology Black start capability Off-grid operability Response time Resilience Potential Summary
Hydrogen production Electrolyzer No Yes (if coupled with renewables) Moderate (minutes) Enables distributed hydrogen production; useful in islanded scenarios but cannot self-start
Hydrogen production Steam methane reforming No No Slow (hours) Suitable for centralized supply with high continuity, but lacks flexibility and emergency use
Hydrogen storage Salt cavern storage No No Moderate (infrastructure-dependent) Suitable for seasonal, large-scale storage; requires energy infrastructure for operation
Hydrogen storage High-pressure tank No Yes Fast (minutes) Modular and deployable; enables local backup supply and fast dispatch during disruptions
Hydrogen transport Hydrogen pipeline No No Fast (seconds-minutes) Enables continuous hydrogen supply, supports regional resilience, but lacks autonomous function
Hydrogen transport HTT No Yes Moderate (minutes-hours) Highly mobile for backup fueling in off-grid or emergency zones
Hydrogen utilization Proton exchange membrane fuel cell Yes Yes Fast (seconds-minutes) Ideal for critical load support, mobile/emergency backup, and black start
Hydrogen utilization Solid oxide fuel cell No (needs heat-up) Yes Slow (1-2 hours) Efficient and stable for long-term off-grid power but slow to start
Hydrogen utilization Hydrogen turbines Yes (some models) Yes (with fuel supply) Moderate (minutes) Suitable for grid-scale backup and restoration, with moderate black start capability
II-C Security Issues of Hydrogen-related Technologies
Although resilience potential can be offered by hydrogen-related technologies in production, storage, transportation, and utilization, their security vulnerabilities also exist in each stage. During the production stage, especially in electrolysis and steam methane reforming, the risk of hydrogen leakage and explosion is significant due to high operating pressures and temperatures. During the storage stage, compressed hydrogen tanks face threats such as high-pressure rupture, while cryogenic liquid hydrogen presents risks of rapid venting. Long-term storage methods like salt caverns must address potential leakage and environmental contamination. During the transportation stage, hydrogen’s small molecular size increases the likelihood of leaks in pipelines or tube trailers, and mobile transport systems are vulnerable to traffic accidents or mechanical damage. Additionally, improper handling during transfer can lead to static ignition or pressure imbalances. During the utilization stage, poor ventilation in confined spaces (e.g., tunnels, vehicles, or buildings) can lead to hydrogen accumulation and fire hazards. Fuel cell systems may also face risks of overheating or malfunction if not properly managed. Across all stages, the growing reliance on automated control and digital monitoring introduces cybersecurity vulnerabilities[59], [60], e.g., hydrogen tank levels and hydrogen demand are vulnerable to false data injection attacks[60]. These security issues highlight the need for rigorous safety standards, advanced monitoring systems, and hydrogen-compatible materials to ensure safe deployment of hydrogen technologies.
TABLE IV: Optimization Challenges Introduced by Adopting Hydrogen in MES for Resilience Enhancement
Index Challenges Hydrogen-related Example Formulas / Model Implications Formulations and Optimization Methods
1 Uncertainties Hydrogen demand varies with mobility or industrial loads $\min_{x\in X}c^{\top}x+\mathbb{E}_{\xi}[Q(x,\xi)]$ Stochastic programming, robust optimization, scenario-based MILP
2 Mixed integer variables Investment locations of electrolyzers and hydrogen tanks $y_{i}^{\text{electrolyzer}}\in\{0,1\},\quad z_{j}^{\text{tank}}\in\{0,1\}$ MILP, MINLP, branch-and-bound, Lagrangian relaxation
3 Multi-timescale variables Slow hydrogen storage level update and fast battery dispatch $H_{t}^{\text{stored}}$ (in hours), $B_{\tau}^{\text{stored}}$ (in 5 minutes) MILP, Nested model predictive control, hierarchical deep reinforcement learning
4 Temporal coupling constraints Hydrogen produced at off-peak hours and used later during peak demand, slow dynamics of hydrogen pipeline pressure $H_{t}^{\text{stored}}=H_{t-1}^{\text{stored}}+H_{t}^{\text{prod}}-H_{t}^{\text{used}}$ MILP, model predictive control, dynamic programming, deep reinforcement learning
5 Spatial coupling constraints Hydrogen produced in region A, stored in region B, and used in region C via pipelines $\sum_{j\in\mathcal{N},j\neq i}h_{ji,t}-\sum_{j\in\mathcal{N},j\neq i}h_{ij,t}=h^{\text{net}}_{i,t},~\forall i\in\mathcal{N}$ MILP, Benders decomposition, distributed optimization
6 Nonlinear & nonconvex constraints Electrolyzer efficiency curve, pressure-flow relationships in pipelines $\eta^{\text{elec}}=f(P_{\text{input}}),\quad Q=k\cdot\sqrt{\Delta P}$ NLP, convex relaxation, piecewise linear approximation
7 Multi-energy coupling constraints Hydrogen enables electricity-gas-heat-transport sector integration $H_{t}^{\text{prod}}=\eta^{\text{elec}}\cdot P_{t}^{\text{elec}}/\omega^{\text{LHV}}$ Integrated MILP/MINLP, Benders decomposition
8 Multi-objective function Tradeoff between cost minimization and resilience/load recovery maximization $\min~\text{Cost}$; $\max~\text{Load~Recovery}$ MILP/MINLP, Multi-objective optimization/reinforcement learning
II-D Optimization Challenges Related to Hydrogen-enabled Planning and Operation Enhancement
Integrating hydrogen into the planning and operation enhancement of MESs introduces a range of complex optimization challenges. Firstly, uncertainty in hydrogen demand complicates the optimization process, often requiring stochastic or robust optimization frameworks. Secondly, many hydrogen-related decisions, including infrastructure investments and on/off operational states of electrolyzers or compressors, involve binary or integer variables, resulting in mixed-integer programming problems. Thirdly, there are multi-timescale decision variables. Due to the physical limits on gas flow, thermal dynamics, and control systems, hourly updates of hydrogen storage level are sufficient for modeling storage dynamics. In contrast, batteries can be dispatched every 5 minutes to capture the fluctuations of renewable generation and load demand. When conducting operation optimization in hours, multi-timescale decision variables are introduced. Fourthly, hydrogen enables cross-temporal operations (e.g., producing and storing energy in low-demand periods for later use), which creates strong inter-temporal coupling and increases the dimensionality of the optimization problem. Fifthly, hydrogen transport via pipelines or trailers introduces spatial coupling, requiring network-based flow models with additional constraints. Sixthly, several key processes (e.g., electrolyzer efficiency, pressure-flow relationships in pipelines, and fuel cell performance) are inherently nonlinear, leading to nonconvex constraints that complicate solution tractability. Seventhly, the integration of hydrogen also deepens the coupling between electricity, gas, heat, and mobility sectors, leading to multi-energy coupling constraints. Finally, there are multi-objective functions in HMESs, e.g., system cost minimization and operation risk minimization, operation cost minimization and load recovery maximization. In summary, the above-mentioned challenges brought by the inclusion of hydrogen call for advanced solution methods, such as mixed-integer nonlinear programming, decomposition techniques, scenario-based planning, robust optimization[61], metaheuristics[60], convex relaxation, model predictive control[62], and deep reinforcement learning[63].
For the ease of understanding, we summarize the optimization challenges introduced by adopting hydrogen in MES for resilience enhancement in Table IV, where $x$ denotes the first-stage decision variables (e.g., hydrogen storage or electrolyzer sizing); $\xi$ represents the uncertainty parameters (e.g., hydrogen demand, renewable availability); $\Xi$ is the uncertainty set; $c^{\top}x$ is the deterministic first-stage investment cost; $Q(x,\xi)$ is the recourse cost in the second stage given scenario $\xi$. $y_{i}^{\text{electrolyzer}}$ is a binary variable indicating whether an electrolyzer is installed at location $i$, $z_{j}^{\text{tank}}$ is a binary variable representing whether a hydrogen storage tank is constructed at node $j$, $y,z=1$ implies the component is built/active, while $y,z=0$ implies it is not. $H_{t}^{\text{stored}}$ and $B_{\tau}^{\text{stored}}$ denote the energy level in hydrogen tank and battery at time $t$ and $\tau$, respectively. $H_{t}^{\text{prod}}$ and $H_{t}^{\text{used}}$ denote the produced hydrogen and the consumed hydrogen at time $t$, respectively. $\mathcal{N}$ is the set of all nodes (e.g., cities, regions, or hydrogen hubs) in the hydrogen network, $\mathcal{T}$ is the set of time periods in the optimization horizon, $h_{ij,t}$ is the amount of hydrogen flow from node $i$ to node $j$ at time $t$, $h_{ji,t}$ is the amount of hydrogen flow from node $j$ to node $i$ at time $t$, $h^{\text{net}}_{i,t}$ is the net hydrogen injection at node $i$ at time $t$, which is positive if hydrogen is injected (e.g., from an electrolyzer) and negative if it is consumed (e.g., by a fuel cell or local demand). $\eta^{\text{elec}}$ is the efficiency of the electrolyzer, $P_{\text{input}}$ is the input electrical power, $f(\cdot)$ represents a nonlinear, possibly nonconvex function of power input, $Q$ is the hydrogen flow rate through a pipeline, $\Delta P$ is the pressure difference across the pipeline section, and $k$ is a physical constant related to pipe geometry and gas properties. $P_{t}^{\text{elec}}$ is the electrical power supplied to the electrolyzer at time $t$, $\omega^{\text{LHV}}$ denotes the lower heating value of hydrogen.
III Hydrogen-enabled Resilience Enhancement Framework for MESs
Figure 2: The framework of hydrogen-enabled resilience enhancement for MESs
As shown in Fig. 2, the framework of hydrogen-enabled resilience enhancement for MESs is provided, where four steps could be identified, i.e., resilience metrics, event-oriented contingency modeling, planning enhancement, and operation enhancement. To be specific, a resilience metric related to the investigated HMESs should be first defined, which would be incorporated into the objectives or constraints of planning or operation problem models. Then, according to the type of extreme events (e.g., hurricane, flood, ice disaster, earthquake, cyber attacks), the corresponding event models could be developed, which consist of event-related physical attributes[64], e.g., movement speed and landfall point of typhoons, and the intensity of earthquakes. Since such attributes depend on some unknown factors, the occurrence of an extreme event is full of uncertainties in the planning stage, e.g., there are many possible landfall scenarios when typhoons are considered[47]. After obtaining the physical attributes of extreme events, the stochastic impact of extreme events on the contingencies of MESs could be developed. Next, taking parameter uncertainty (e.g., renewable generation and multi-type loads), event uncertainty, and event-oriented contingency uncertainty into consideration, scenario generation could be conducted in preparation for planning and operation problems. In the planning stage, some permanent facilities will be deployed, e.g., SHSs, HESSs, hydrogen pipelines, and HRSs. During the operation stage, several kinds of responses can be made, e.g., preventive response, emergency response, and restoration response. During the preventive response, measures such as MHER location allocation, hydrogen procurement, and HESS pre-charging can be taken. During the emergency response, some measures are taken to minimize load shedding, including MHER scheduling, natural gas pipeline-assisted hydrogen use, HESS discharging, hydrogen demand adjustment, and hydrogen transformation. In the restoration stage, mobile hydrogen generation and storage, MHER routing and scheduling, HESS scheduling, hydrogen fuel cell scheduling, and hydrogen pipeline scheduling could be adopted to restore loads. By evaluating the practical system performance in the aspect of resilience, the rationality of the developed resilience metrics and event-oriented contingency modeling could be further improved.
IV The Definitions and Metrics of HMES Resilience
Before enhancing the resilience of an HMES, its definitions and metrics should be established, since resilience metrics will be incorporated into the objectives or constraints of planning/operation problem models. In the following parts, HMES definitions and architectures are introduced. Then, the definitions of HMES resilience are described. Next, typical resilience metrics in existing works are classified. Finally, a summary of the performance metrics is provided.
IV-A HMES Definitions and Architectures
Similar to the MES concept[65], an HMES is a system in which hydrogen, electricity, heat, cooling, gas, and transport interact optimally with each other at various levels to increase technical, economic, and environmental performance relative to “classical” energy systems whose sectors are treated separately. According to the spatial perspectives, HMESs can be divided into building-level, district-level, region-level, and nation-level. For simplicity, two typical architectures of HMESs are introduced, i.e., building-level HMESs and region-level HMESs. As shown in Fig. 3, the framework of a zero-carbon building-level HMES can be observed, and the corresponding practical system picture can be found in [66]. In Fig. 3, there are four kinds of energy flows, i.e., electricity flow, heat flow, cooling flow, and hydrogen flow. As far as electricity flow is concerned, the electricity generated from photovoltaic (PV) panels and the fuel cell is used to satisfy the electricity demand of an electric boiler, a ground-source heat pump (GSHP), an absorption chiller, an electrolyzer, and electric demand in a commercial building. As to heat flow, the heat demand is satisfied by the electric boiler, GSHP, fuel cell, and hot water tank. Similarly, cooling demand is satisfied by the GSHP, absorption chiller, and cold water tank in the cooling flow. As to hydrogen flow, PV generation can be used for the electrolyzer to produce hydrogen, which will be stored in the hydrogen tank for future use by driving fuel cells. When the hydrogen stored in the hydrogen tank is not enough, hydrogen purchasing from the market is required. In Fig. 4, the architecture of a region-level HMES can be observed[67], which consists of three types of communities, e.g., residential communities, commercial communities, and industrial communities. In residential and commercial communities, electricity/cold/heat/hydrogen demands can be identified. In industrial communities, electricity/hydrogen demands can be observed. All communities are connected through electricity transportation networks, hydrogen transmission pipelines, and transportation networks. With the help of transportation networks, MHERs (e.g., HTTs and HFCVs can be redistributed among multiple communities for improving network resilience, e.g., providing emergent power supply or hydrogen supply to a community when it suffers from an extreme event (e.g., power outages or hydrogen pipeline leakage).
Figure 3: Typical HMES architecture (building-level)
Figure 4: Typical HMES architecture (region-level)
IV-B Definitions of HMES Resilience
Figure 5: System performance curves related to MES and HMES operation
Before describing the definitions of HMES resilience, we first provide two system performance curves related to MES and HMES under extreme events in Fig. 5, respectively. Although the typical curve under MES cannot perfectly capture the realistic system performance trajectories (e.g., local performance fluctuations) as illustrated in [68, 69, 70], it can reflect the system’s ability to prepare for, absorb, and recover from disturbances and has been widely used in existing works[50],[71, 72, 73, 74, 75],[115]. In addition, While it is not tailored to hydrogen-enabled MESs, the performance trajectory can be applied to hydrogen-enabled MESs after minor modifications. By comparing the differences between two curves, the physical and functional features of hydrogen can be observed. Firstly, we introduce the whole process of resilience enhancement under MES. In existing works, load loss is a widely used metric of system performance. Under this premise, the curve under MES consists of normal operation stages and three kinds of responses (including pre-event preventive response, during-event emergency response, and post-event restoration response). In the normal operation stage, the optimal economic operation is typically conducted from time $t_{0}$ to time $t_{1}$. After detecting the extreme event information, the preventive response will be carried out from time $t_{1}$ to time $t_{2}$, e.g., purchasing mobile electric vehicle resources. During the extreme event, the emergency response is adopted to minimize the load shedding from time $t_{2}$ to time $t_{3}$ so that the extent of system performance degradation could be alleviated (note that the system performance can be maintained for a short time due to the deployed mobile resources in the preventive response). Then, the degraded state could be observed from time $t_{3}$ to time $t_{4}$ at the end of event. From time $t_{4}$ to time $t_{5}$, the restoration response is achieved to restore critical loads [74] like hospitals and water pumps by adopting temporary measures, such as grid-island operation using remaining storage and mobile energy storage units (e.g., electric vehicles). These temporary measures restore partial or prioritized functionality, even if the underlying infrastructure (e.g., transmission lines and distribution equipment) is not yet fully repaired. After stabilizing essential services, the system proceeds to recover full infrastructure, e.g., repairing line and equipment.
Similarly, after integrating hydrogen into MESs, the system performance curve should reflect both the physical and functional characteristics of hydrogen. In the preventive stage, additional mobile energy resources (e.g., HTTs for gas delivery and liquid-hydrogen tankers, as well as HFCVs/HFCBs used as mobile generators) can be deployed, so that system performance can be sustained longer in the emergency period. During the emergency response, if a grid outage coincides with a hydrogen-pipeline leak, part of critical demand can be supplied by fuel cells. Owing to pipeline linepack and pressure regulation, fuel cell stack-inlet pressure/flow conditions can be temporarily maintained, allowing the fuel cell to sustain its set power and thereby reduce load shedding as long as the available upstream pressure/flow remains above the requirements of hydrogen regulator and fuel cell stack. If the leak reduces the available supply below those requirements, hydrogen starvation occurs[116], leading to a rapid voltage drop and possible protective shutdown, which increases load shedding. Although linepack buffering may delay this adverse impact, fuel-cell power loss can worsen even after the initiating event ends once the buffer is depleted. Since more resources are available for deployment, an HMES can exhibit a higher performance level in the degraded state, and an example of this phenomenon can be seen in Fig. 7 of [117]. Utilizing mobile resources or conducting pipeline reconfiguration, the recovery trajectory of HMES can also be raised. Compared with MESs, hydrogen-related equipment failure in HMESs may require longer restoration time as shown in Fig. 5, which rely on random transportation network state and hydrogen flow dynamics as shown in Fig. 6 of [120] in the revised manuscript.
Based on the above description, HMES resilience can be defined as the capacity of an HMES to withstand the disturbance from an extreme event (i.e., event progress during $t_{2}$ and $t_{3}$), and to rebuild and renew it afterwards[9]. Similarly, HMES resilience can also be described as the capacity to limit the system performance degradation and the duration of the degraded state in order to maintain critical services following an extreme event[5]. In addition to energy resilience[118], carbon resilience is also proposed in existing works, which describes an energy system’s ability to maintain or recover to its current and predicted emission level during and after shocks such as renewable intermittency, public health emergencies, or climate-related disruptions[53]. Although many definitions related to HMES resilience are proposed, they are originally borrowed from power grids and MESs, which can not truly reflect the physical and functional features of hydrogen[18][28][71, 72, 73, 74, 75]. For example, when an extreme event (e.g., ice disaster) causes simultaneous power outages and hydrogen pipeline leakage, even if no load loss is incurred due to backup sources, hydrogen fuel cells can not operate normally due to the loss of pressurized hydrogen supply. At this time, no resilience enhancement measure is required since there is no system performance degradation (i.e., load loss). In fact, hydrogen-related systems need immediate repair so that further pressure reduction can be avoided. Otherwise, once backup sources are unavailable, the whole HMES may collapse and result in larger negative social-economical impacts. Therefore, the definition of HMES resilience should incorporate any negative impacts (e.g., hydrogen availability and hydrogen-demand loss) of extreme events on hydrogen energy resources into consideration.
IV-C Metrics of HMES Resilience
In existing works, resilience metrics related to HMESs can be divided into three types, i.e., time indexes, load loss indexes, and load rate indexes[10]. To be specific, time indexes are used to describe the duration of each operation stage, while load loss indexes are adopted to describe the extent of load shedding in different stages. Smaller time and load-loss indices are preferred. Load rate indexes are used to evaluate the rates of load loss or load recovery. Thus, a smaller load loss rate and a larger load recovery rate are expected. In existing works related to HMES resilience, the most widely used resilience metrics for HMES operation are summarized in Table V, and detailed descriptions are given below.
TABLE V: Summary of widely-used resilience metrics for HMES operation
Metrics Definitions References
LSR Load served ratio [71], [77], [50]
ELNS Expected load not served [29], [72, 73, 76], [78]
ERL Expected restored load [74], [79, 80, 81]
WSLL Weighted sum of load loss [47], [18], [82, 83, 84]
$L$ Load loss [31], [75], [85, 50, 86, 87, 11, 88, 89]
CRI Comprehensive resilience index [28]
IV-C1 LSR
The load served ratio (LSR) is adopted to describe the operation resilience of hydrogen-powered microgrids[71], which is defined by
$\textbf{LSR}=\sum\nolimits_{i\in\mathbf{N}}\lambda_{i,t}P_{i,t}/\sum\nolimits_{i\in N}P_{i,t},$ (1)
where $\mathbf{N}$ denotes the set of buses, $\lambda_{i,t}$ is a binary variable to indicate whether the load at bus $i$ at slot $t$ is served or not. When the corresponding bus is served, the value of $\lambda_{i,t}$ is one. Otherwise, its value is zero.
IV-C2 ELNS
In [72] and [73], Cai et al. and Jordehi et al. proposed the same resilience metric in the preventive stage for hydrogen fuel station-integrated power systems and industrial energy hubs with electric, thermal and hydrogen demands, respectively. The resilience metric is the expected load not served (ELNS), which can be described by
$\textbf{ELNS}=\mathbb{E}_{s}\{\sum\limits_{i}\sum\limits_{t}(w_{e}P_{i,t,s}^{\text{shed}}+w_{h}H_{i,t,s}^{\text{shed}}+w_{g}G_{i,t,s}^{\text{shed}})\},$ (2)
where $s$ denotes the index of uncertainty scenarios; $t$ denotes the time slot index; $i$ denotes the bus index; $w_{e}$, $w_{h}$, and $w_{g}$ denote the weight coefficients related to electric load shedding, hydrogen load shedding, and heat load shedding, respectively; $P_{i,t,s}^{\text{shed}}$, $H_{i,t,s}^{\text{shed}}$, and $G_{i,t,s}^{\text{shed}}$ denote the quantity of electric, hydrogen, and heat load shedding, respectively. Similarly, expected thermal load not served and expected electric load not served are considered for hydrogen-penetrated distribution systems in the restoration response stage[76].
IV-C3 ERL
In the restoration response stage, expected restored load (ERL) maximization was considered for hydrogen-based microgrids in [74], which is given by
$\textbf{ERL}=\sum\nolimits_{\theta}\varrho_{\theta}\Phi_{\theta},$ (3)
where $\theta$ denotes the uncertainty scenario index, $\varrho_{\theta}$ denotes the scenario probability, $\Phi_{\theta}$ denotes the served weighted load of heating, cooling, electricity, gas, and hydrogen demands[79, 80, 81].
IV-C4 WSLL
In [18], Li et al. proposed a resilience metric in the emergency response stage for a hydrogen-penetrated multi-energy supply microgrid, which is the weighted sum of load loss (WSLL) related to electricity, heat, and gas, i.e.,
$\textbf{WSLL}=\sum\nolimits_{t}(\gamma L_{t}^{e}+\delta L_{t}^{h}+\epsilon L_{t}^{g}),$ (4)
where $\gamma$, $\delta$, and $\epsilon$ denote the weights related to electric load shedding, heat load shedding, and gas load shedding, respectively. $L_{t}^{e}$, $L_{t}^{h}$, and $L_{t}^{g}$ are the quantities of electric load shedding, heat load shedding, and gas load shedding, respectively. When just electric load shedding is involved, WSLL is reduced to load loss (LL), which was adopted in many works[75].
IV-C5 CRI
Note that the above indexes mainly focus on load loss, restored load, or served load ratio, they neglect the temporal performance that load can be restored. To overcome the challenge, Zhao et al. proposed a comprehensive resilience index (CRI) to reflect the temporal performance and system overall performance[28], i.e.,
$\textbf{CRI}=\beta R_{0}(t_{5})+(1-\beta)R_{1}(t_{5}),$ (5)
$R_{\sigma}(t_{5})=\frac{\int_{t_{2}}^{t_{5}}F(t_{0})^{\sigma}[F(t)>F^{\min}]dt}{\int_{t_{2}}^{t_{5}}F(t_{0})^{\sigma}dt},$ (6)
where $t_{2}$ and $t_{5}$ denote the event happening time and full recovery time as shown in Fig. 4, $F^{\min}$ denotes the minimum system performance requirement, and $F(t)>F^{\min}]=1$ if $F(t)>F^{\min}$. Otherwise, $F(t)>F^{\min}]=0$. $\sigma\in\{0,1\}$. Thus, $R_{0}(t_{5})$ can reflect the time ratio that can satisfy the minimum system performance requirement, and $R_{1}(t_{5})$ can reflect the overall service performance.
Summary: In existing works, many load-based metrics have been widely adopted, which often translate into economic damage or loss. However, they often overlook the social impacts of service disruption. In practice, the number and types of customers affected are also critical indicators of resilience, especially during extreme events [90],[91]. For instance, shedding the same amount of load for a hospital and a residential building may lead to totally different consequences. The reason is that the former relies on power for life-saving operations, while the latter may experience only reduced comfort. Therefore, to develop a comprehensive resilience metric, it is necessary to combine load-based and customer-based metrics, as mentioned in [92]. In addition, existing metrics can not reflect the degraded state (e.g., HESS failure or hydrogen pipeline leaking) of HERs during an extreme event. Therefore, more comprehensive metrics are expected.
V Event-oriented Contingency Modeling in HMESs
Different high-impact and low-probability events would incur different types of contingencies. According to the stochasticity and contingency type, existing contingency modeling could be classified into four types as shown in Fig. 6, i.e., deterministic single-type contingencies[28, 29, 30],[75],[50],[79],[80],[83],[84],[93],[94], deterministic multi-type contingencies[103], stochastic single-type contingencies[47],[11],[22],[23],[72],[73],[78],[77],[87],[95, 96, 97],[98], and stochastic multi-type contingencies[99],[100],[46],[18]. Since deterministic contingencies can not represent the uncertain characteristics of high-impact and low-probability events, this paper mainly focuses on stochastic contingencies.
Figure 6: The classification of event-oriented contingency modeling
V-A Stochastic Single-type Contingency
V-A1 Main grid disconnection model
There are two types of main grid disconnection models, i.e., the $T-\tau$ criterion emergency islanding model and the stochastic grid outage model. The former is widely used for the microgrid islanding model, which means that a microgrid can operate in an islanding mode for continuous $\tau$ periods that could occur at any period of scheduling horizon $T$, which can be described by[95]
$y_{h,t}=0,~\forall~t\in~[h:\min(T,h+\tau-1)],$~ (7)
$\sum\nolimits_{t\in[\gamma:T]}y_{h,t}=T-\gamma-\tau+1,h=\gamma,$ (8)
where $\gamma$ denotes the current scheduling slot, $y_{h,t}$ denotes the connection state between a microgrid and the main grid at slot $t$ under the emergency islanding scenario $h$. $h=r$ denotes the scenario $h$ with starting disconnection slot $\gamma$. When $y_{h,t}=1$, it means that the microgrid is connected to the main grid at slot $t$ under the scenario $h$. Otherwise, $y_{h,t}=0$. The stochastic grid outage model is used to generate main grid power outage scenarios with different starting times and ending times according to a given random distribution[78]. Compared with the $T-\tau$ criterion emergency islanding model, the stochastic grid outage model can generate power outage profiles with different durations, instead of a constant duration.
V-A2 Transmission line fault model
The transmission line failure model is used to generate a large number of line failure scenarios, where each scenario consists of combinations of all random line states[96]. To be specific, when the randomly generated values are smaller than the corresponding line failure probability, the considered line is broken. Otherwise, the considered line is still in a normal state. As shown in [97], the failure probability of line $z$ can be described by
$P_{k}=1-(1-P_{0})^{l_{z}},$ (9)
where $P_{0}$ denotes the failure probability of unit length transmission line and $l_{z}$ denotes the length of line $z$. Note that $P_{0}$ is the function of attributes of ice disasters, e.g., ice load and wind speed. In addition, the NkN-k line contingency model is adopted in many existing works. To be specific, let $v_{i,j}$ be the connection state of line $(i,j)$, which is a binary variable. Moreover, $v_{i,j}=1$ if the line $(i,j)$ operates under normal state, and $v_{i,j}=0$ if it is damaged. Then, we have $\sum\nolimits_{(i,j)\in\mathcal{L}}(1-v_{i,j})\leq k$. Note that such a model is considered in many existing works[11],[77].
When the starting period is considered, the NkN-k contingency model can be reformulated as follows, i.e.,
$\sum\nolimits_{(i,j)\in\mathcal{L}}(1-v_{i,j})\leq k,~\sum\nolimits_{t\in\mathbf{T}}\varepsilon_{t}\leq 1,$ (10)
where $\varepsilon_{t}=1$ means that the contingency begins at period $t$, and $\mathbf{T}=\{1,2,\cdots,24\}$[22].
Since many extreme events have spatial features (e.g., typhoons), a location-aware line outage uncertainty model can be adopted, e.g., a line outage uncertainty set under a typhoon scenario $s$ can be defined by[47]
$\mathbb{U}_{s}=\Big{\{}\boldsymbol{\vartheta}_{s}|K_{s}\geq\sum\limits_{\{i,j\}\in\Omega_{\text{PLrisk}}}(\nu_{i,j}-\vartheta_{i,j})\Big{\}},~\forall~s\in\boldsymbol{S},$ (11)
where $\nu_{i,j}=1$ if $\{i,j\}\in\Omega_{\text{PLrisk}}$; $\vartheta_{i,j}=0$ indicating the line outage, while $\vartheta_{i,j}=1$ means that there is no line outage; $K_{s}$ denotes the allowable maximum number of line outages; $\boldsymbol{\vartheta}_{s}=(\vartheta_{i,j})|_{\{i,j\}\in\Omega_{\text{PLrisk}}}$.
V-A3 Hydrogen pipeline fault model
According to [111], the dynamic probability model related to hydrogen pipeline faults can be described by a lognormal cumulative distribution function $\psi(\cdot)$ as follows,
$\text{Pr}_{m,n,k,t}=\psi\Bigg{(}\frac{\ln(\zeta_{m,n,k}\sum\nolimits_{\tau\leq t}r_{m,n,k,\tau})}{\chi_{m,n,k}}\Bigg{)},$ (12)
where $r_{m,n,k,\tau}$ denotes the rainfalls suffered by the segment $k$ of hydrogen pipeline $mn$ at slot $\tau$, $\zeta_{m,n,k}$ and $\chi_{m,n,k}$ are two parameters estimated from empirical data. Then, the failure probability of the hydrogen pipeline $mn$ at slot $t$ is give by
$\text{Pr}_{m,n,t}=1-\Pi_{k=1}^{N_{s}}(1-\text{Pr}_{m,n,k,t}),$ (13)
where $N_{s}$ is the number of segments in hydrogen pipeline $mn$.
V-A4 Single-type device impact model
Device fragility models related to poles and towers, insulators, transformers, and photovoltaic panels under extreme weather can be calculated according to the related physical parameters. More details could be found in [23],[87],[98]. For example, photovoltaic panel icing has a large impact on the intensity of insolation, which can lead to a decrease in photovoltaic generation power. Let $P_{t}^{\text{fault}}$ be the maximum generation output when panels experience freezing rain disaster at slot $t$. Then, we have[23]
$P_{t}^{\text{fault}}=\Gamma\Big{(}G\exp\big{(}-\frac{1.5\rho_{\text{frz}}H_{t}^{\text{PV}}}{\rho_{\text{ic}}r_{\text{ef}}}\big{)}\Big{)},$ (14)
where $\Gamma(\cdot)$ denotes a function; $G$ is the insolation intensity on the surface of the ice cover; $\rho_{\text{frz}}$ and $\rho_{\text{ic}}$ are the intensities of freezing rain and ice, respectively; $r_{\text{ef}}$ denotes the effective grain radius; $H_{t}^{\text{PV}}$ denotes the ice thickness on the panel surface.
V-A5 Repairable equipment failure model
Since some devices caused by extreme events (e.g., hurricanes) are repairable, a Markov two-state transition model was adopted to describe the “operation-failure-operation” cycle process of wind turbine or photovoltaic equipment in [101]. To be specific, the transition from normal operation state to the failure state in a continuous-time Markov process happens according to a transition rate $\lambda_{o}$. Similarly, the transition from the failure state to normal operation state happens according to a transition rate $\lambda_{f}$. Let $T_{otf}$ and $T_{fto}$ denotes the average sojourn time of operation state and failure state, respectively. Then, $T_{otf}=1/\lambda_{o}$ and $T_{fto}=1/\lambda_{f}$, which also represent the mean time to failure and mean time to repair, respectively. Moreover, the above-mentioned sojourn time follows an exponential distribution.
V-B Stochastic Multi-type Contingency
V-B1 Multi-device fault model
Multi-device fault model was adopted in [46] to capture the uncertain contingencies of electrolyzer units, fuel cells, and gas turbine units caused by disastrous events. Let $N_{g}$ be the minimum number of components under normal operation. Then, the multi-device fault model can be described by an uncertainty contingency set as follows,
$\Omega=\{\boldsymbol{u}|\sum\nolimits_{e}u_{\text{EL},e}+\sum\nolimits_{f}u_{\text{FC},f}+u_{\text{GT}}\geq N_{g}\},$ (15)
where $\boldsymbol{u}=u_{\text{EL},e},~u_{\text{FC},e},~u_{\text{GT}}$. Moreover, $u_{\text{EL},e}$$u_{\text{FC},e}$$u_{\text{GT}}$ are binary variables related to electrolyzers, fuel cells, and a gas turbine, respectively. When there is a fault in a component, the corresponding value is zero. Otherwise, its value is one. Similar to the above multi-device fault model, NkN-k contingency model was considered in [99] for the stochastic failures of critical components (e.g., fuel cell stacks).
V-B2 Temporal-spatial destructive model
In [18], a temporal-spatial destructive model was adopted to describe the contingencies of electricity/gas/heat networks under the impact of the hurricane, which is varying and highly related to spatial locations of components in electricity/gas/heat networks. Firstly, the grid division method is used to describe the damage of hurricanes at slot $t$. Specifically, a three-dimensional matrix $G_{\text{state}}^{r,c,t}$ is constructed and its element is equal to one if the rectangle area with location $(r,c)$ belongs to the hurricane area at slot $t$. Otherwise, its element is equal to zero. Secondly, a two-dimensional matrix $G_{\text{loc}}^{r,c}$ is created to describe which components belong to the rectangle area $(r,c)$. Based on the above matrixes, the damaged level could be obtained. Continually, a node export ability in the network can be calculated by the percentage of survival paths from all energy sources to the current node.
Summary: Compared with stochastic single-type contingencies, stochastic multi-type contingency modeling has a stronger representation ability. Even so, the multi-device fault model still cannot capture the dynamics of contingencies related to extreme events, since the sequences and starting times of fault appearances are not considered. In fact, extreme events (e.g., hurricanes and floods) have temporal and spatial attributes, which may lead to multi-type contingencies with different sequences and starting times. In addition, some cascading contingencies incurred by cyber-physical attacks are not considered in existing works[102]. Above all, existing works rarely consider hydrogen-specific vulnerabilities. In fact, there are many physical vulnerabilities and cyber vulnerabilities in HMESs. In physical aspects, hydrogen tends to cause material embrittlement, particularly in metals such as steel, can lead to long-term degradation of pipelines, storage tanks, and connectors. These structural vulnerabilities increase the likelihood of mechanical failure under normal operations or stress conditions such as temperature fluctuations and pressure surges during extreme events. In cyber aspects, a hacker manipulating the pressure control in a hydrogen storage tank could trigger an overpressure event, potentially causing mechanical damage or explosion. Similarly, false data injection into hydrogen demand forecasts or electrolyzer schedules could destabilize coordinated operations across the electricity-hydrogen interface[60], leading to misallocation of resources or even blackouts. However, most contingency models treat hydrogen assets as idealized components with fixed performance characteristics in existing works, ignoring the risks posed by leakage, material fatigue, and cyber intrusion. Therefore, future research should develop risk-aware, vulnerability-inclusive contingency modeling that capture both the physical fragilities and cyber risks related to hydrogen.
VI Hydrogen-enabled Planning Enhancement for MESs
In the stage of planning, many hydrogen resources will be deployed permanently to enhance the HMES resilience, such as SHSs, HESSs, hydrogen pipelines, HRSs, fuel-cell-driven CCHP, and hydrogen-related production facilities. According to the types of investment-related hydrogen resources, existing works can be generally classified into two types, i.e., single-type hydrogen facilities and multi-type hydrogen facilities. In the following parts, related works will be introduced in detail.
VI-A Single-type Hydrogen Facilities
Many resilience-oriented planning methods have been developed to decide the capacity, location, and quantity of single-type hydrogen facilities (e.g., HESSs, HRSs, and H2Ps). For example, Yan et al.[103] proposed a resilient planning method for power systems to decide electric ESSs and HESS capacities against heatwave events. Firstly, the impact model of heatwave events on the power demand and photovoltaic generation was developed. Then, a total cost minimization problem was formulated with the consideration of the daily investment cost, operation/maintenance cost, fuel cost, power curtailment penalty cost, and so on. Finally, CPLEX was adopted to solve the formulated problem. The proposed planning method can combine the advantages of electric ESSs and HESSs, i.e., HESSs are adopted for long-term storage before heatwave events happen, whereas electric ESSs are responsible for short-term intra-day power balance. During the heatwave events, HESS adopted a discharging strategy to relieve the negative impact of heatwave events. Similarly, Zhao et al.[28] investigated a resilient planning problem to decide the photovoltaic panels, HESS, and electric ESS capacities for a hydrogen-integrated airport energy system considering main grid power outages incurred by extreme events. To achieve the above aim, a planning method was proposed to minimize the total economic cost that consists of the capital cost, operational cost, and carbon emission cost. Simulation results showed that the comprehensive resilience index can increase by 18% when the hydrogen-penetration ratio in the airport energy system increases from 20% to 40%. In the above works, the deterministic effects of extreme events were considered. In practice, those effects may be stochastic. For example, Gao et al.[96] proposed a two-stage resilience planning method for an integrated electricity-gas energy system with the consideration of stochastic line fault impacts incurred by typhoons. As shown in Fig. 7, the first stage solves the investment decision problem to decide the locations and capacities of HRSs and gas-fired distributed generators under the given quantity constraints. The second stage solves the system operation optimization problem in all fault scenarios, which were generated based on the Monte Carlo method and line fragile probability models under typhoons. The decision variables are upstream purchase, electric/gas load shedding, fuel cell generation, and gas-fired distributed generation. Finally, a single-level MISOCP problem was formulated to minimize the sum of investment cost and expected operation cost and solved by the Gurobi solver. Simulation results showed that the proposed planning method can reduce the total cost compared to other four baselines by 6.36%-50.54%, which focus on only resilience (load shedding), only investment economics, only HRS planning, and only gas-fired distributed generators.
Figure 7: The framework of two-stage resilience planning method
TABLE VI: Summary of Planning Enhancement for HMES (Single-type Facility)
Ref. Object Events and contingencies Hydrogen facility investment decisions Demand types Parameter/event uncertainty Optimization objectives Problem types, solving methods
[103] Power systems based on HESSs Deterministic impacts caused by heatwave events HESS (capacity) Electricity $\times$ Minimizing the total cost Single-level deterministic MILP, CPLEX
[28] Hydrogen-integrated airport energy system Insufficient power supply capacity caused by extreme events HESS (capacity) Electricity, hydrogen $\times$ Minimizing the overall economic cost Single-level deterministic MILP, CPLEX
[96] Integrated electricity-gas energy system with HRSs Stochastic line outages incurred by typhoons HRS (capacity, location, quantity) Electricity, gas $\times$ Minimizing the sum of investment cost and expected operation cost, minimizing operation costs under scenarios Two-stage MISOCP, Gurobi
[104] Hydrogen microgrids Deterministic line outages by natural disasters HESS (capacity, location, quantity) Electricity $\times$ Minimize the total cost; maximize the operation cost by selecting the worst extreme scenario; minimize load shedding and switch action costs Tri-level robust MILP, mixed-integer L-shaped method
[46] Electricity-hydrogen island energy system Stochastic multi-device failures and power interruptions during extreme events HESS (capacity) Electricity, hydrogen Wind power and electric load Maximizing the total profit, minimum value in the third level, maximizing profit minus load shedding cost Tri-level robust MILP, nested C&CG
[106] European electricity systems Low renewable generation during Dunkelflaute events HESS (location, capacity) Electricity Renewable generation Minimize total annualized costs; maximizes the operation cost by selecting the worst scenario; minimize the operation cost after the uncertainty realization Tri-level robust MILP; C&CG
[47] Electricity-hydrogen integrated energy system Stochastic line outages incurred by typhoons H2P (capacity, location, quantity) Electricity, hydrogen Event uncertainty Minimizing investment cost and expected worst-case system cost Tri-level stochastic robust optimization problem, nested C&CG with progressive hedging
[107] Gas-electric-hydrogen microgrids NkN-k contingencies caused by extreme events HESS (location, quantity) Electricity, heat $\times$ Minimizing the total cost, maximizing load restoration Bi-level MILP, CPLEX
Figure 8: The framework of the tri-level robust planning method
In [104], Qu et al. investigated a tri-level robust planning problem for enhancing the resilience of distribution networks considering deterministic line faults. The optimization objective is to minimize the weighted sum of annual investment cost and annual load shedding cost, and decision variables of the first-level problem are the locations, quantities, and capacities of hydrogen microgrids and soft open points. To generate the stochastic line fault scenarios, a multi-type natural disaster model was developed. To solve the formulated tri-level MILP problem, a mixed-integer L-shaped method was proposed. Simulation results indicated that cooperative planning of hydrogen microgrids and soft open points can significantly enhance the resilience of the distribution network and reduce load shedding by 68.4%. In [46], Gu et al. proposed a tri-level planning method to determine the capacities of HESSs and gas turbines for an electricity-hydrogen island energy system while taking the uncertainties of wind power, electric load, and line faults into consideration. As shown in Fig. 8, three levels could be identified. Note that levels 2 and 3 intend to maximize the profit under the worst-case scenario, while level 1 intends to maximize the sum of the maximum profit obtained from level 2 minus the investment cost, which consists of the annual average investment cost, the annual average depreciation cost, and the annual average fixed operation and maintenance cost. To achieve the above aim, a $\max$-$\min$-$\max$ tri-level problem was formulated under different representative days. Finally, a nested C&CG algorithm was adopted to solve the formulated problem. Compared with the case that only considers worst-case contingency, the proposed planning method can reduce load shedding by 86%. Similarly, Bernecker et al.[106] investigated a capacity planning problem for fully decarbonized European electricity systems based on tri-level robust optimization framework considering Dunkelflaute extreme weather events (i.e., prolonged periods of low wind and solar availability). The first level focuses on minimizing investment cost, the second level focuses on maximizing the total system cost by finding the worst-case realizations, and third level focus on minimizing the total system cost under the given uncertainty realization. By merging the dual problem of the third-level subproblem with the second-level subproblem, an equivalent maximization problem can be derived. Finally, the $\min-\max$ problem could be solved by C&CG algorithm. Simulation results indicated the importance of considering widespread regional weather events when designing a robust Europe electricity system.
Although stochastic effects of an event have been considered in the above-mentioned works, they neglect the event uncertainty. For example, when typhoons are considered in the resilient planning of an MHES, there are many possible scenarios related to typhoon spatial distributions, and stochastic effects under each scenario are incurred. As shown in Fig. 9, line outages within the high-risk areas (e.g., high wind speed areas) under a typhoon scenario may happen. By taking event uncertainty into consideration, Wen et al. proposed a tri-level planning method for a hydrogen-electricity integrated energy system[47]. To be specific, a Wasserstein generative adversarial network with a gradient penalty-based method was presented to generate the typical typhoon scenarios. Then, spectral clustering was adopted to reduce typhoon scenarios. Next, a tri-level optimization problem was formulated, where the upper level intends to decide the investment decisions related to power line hardening and H2P facilities, the middle level intends to find the worst-case line outage event under each typhoon scenario, while the lower level intends to minimize the weighted sum of load shedding related to power demand and hydrogen demand under each line outage event. The above tri-level problem was finally solved by nested C&CG with a progressive hedging algorithm. Simulation results indicated that the proposed planning method can reduce load shedding by 95.7% compared with existing uncertain planning methods.
Different from above-mentioned works that deal with uncertainty in optimization, leader-follower structure is adopted for modeling the planning problem with the consideration of contingencies. To be specific, Oh et al.[107] proposed a two-level resilient planning approach for gas-electric-based multi-energy systems, including a microgrid, when NkN-k contingencies are considered. The upper level intends to minimize the total cost of the microgrid by deciding the capacity, location, and quantity of each component (e.g., distributed generators, electric ESSs, HESSs, wind turbines, photovoltaic panels). The lower level intends to maximize the restored load. Simulation results showed that the proposed approach can enhance the system’s resilience by the adoption of the microgrid and reduce expected electric load loss by 30%-34%.
Figure 9: Illustration of line outages in an electricity-hydrogen integrated energy system under typhoons
TABLE VII: Summary of Planning Enhancement for HMES (Multi-type facilities)
Ref. Object Events and contingencies Hydrogen facility investment decisions Demand types Parameter/event uncertainty Optimization objectives Problem types, solving methods
[108] Integrated gas and electricity systems with HESSs Two kinds of deterministic extreme weather events HESS, hydrogen production facilities (capacity, location) Electricity, heating $\times$ Minimizing the total cost Deterministic single-level MINLP, Xpress solver
[99] Hydrogen-based microgrids NkN-k stochastic device failure caused by disastrous events HESS, hydrogen-driven CCHP, SHS (capacity, location, quantity) Electricity, heating, cooling Renewable generation and demand Minimizing the total cost, minimizing the operation cost Two-stage stochastic MINLP, dual cutting-plane decomposition algorithm
[109] Hydrogen-based seaport MESs NkN-k line outage contingencies Fuel cell electric truck, HRS, renewable generator, and remote control control (capacity, location) Electricity, heating, cooling Renewable generation, demand, and price Minimizing investment cost and expected operation cost Two-stage stochastic MINLP, Gurobi
[110] Hydrogen-enriched power distribution networks NkN-k contingencies caused by extreme events HESS, MHER (location, capacity) Electricity, hydrogen Renewable generation, electrical and hydrogen load, and network contingencies Minimize the total cost; minimize the expected operation cost Risk-averse two-stage stochastic MILP, Gurobi
[22] Hydrogen-based microgrids and hydrogen fuel cell bus NkN-k line contingencies incurred by extreme weather HRS and hydrogen fuel cell electric bus (capacity, location, quantity) Electricity, hydrogen $\times$ Maximizing profit, minimum value of maximized load restoration, maximizing load restoration Tri-level MILP, nested C&CG
[111] Electricity-hydrogen distribution networks Stochastic power line and hydrogen pipeline contingencies caused by extreme weather Hydrogen pipeline hardening, HESS (capacity) Electricity, hydrogen $\times$ Minimizing the maximum load shedding under the worst-case scenario, minimizing load shedding Distributionally robust MISOCP, C&CG algorithm
[112] Hydrogen-electrical microgrids for sea islands NkN-k contingencies caused by extreme wind events HESS (location, capacity), hydrogen vessel purchasing and routing Electricity Electricity demand, renewable generation, line outages, hydrogen vessel availability Minimize total annualized costs; minimize operational and maintenance costs under the worst-case probability distribution Distributionally robust MILP, C&CG with strong cutting planes
[113] Renewable-dominated power systems Stochastic line and tower damage under extreme weather Fuel cell, battery, electrolyzer, and hydrogen tank (capacity, rated power, location) Electricity, hydrogen Renewable generation and demand Minimizing configuration cost and operation cost Bi-level problem, PSO and MILP
TABLE VIII: Modeling Frameworks of Planning Enhancement Problem for HMESs
Modeling framework Model explanation Mathematical formulation Solution method
Single-level deterministic optimization Classical optimization with known parameters; no uncertainty or hierarchical decision structure $\min_{x\in X}c^{\top}x+f(x)$ Linear/Nonlinear/Mixed-integer programming solvers (e.g., Gurobi, CPLEX, IPOPT, Xpress)[103][108][28]
Two-stage stochastic programming First-stage “here-and-now” decisions made before uncertainty is realized; second-stage “wait-and-see” recourse decisions after uncertainty is revealed $\min_{x\in X}c^{\top}x+\mathbb{E}_{\xi}[Q(x,\xi)]$, where $Q(x,\xi)=\min_{y}\{q^{\top}y:Wy+Tx\geq h\}$ Column-and-Constraint Generation (C&CG)[111], dual cutting-plane decomposition[99], commercial solvers (e.g., Gurobi[109][110][96])
Tri-level robust optimization Hierarchical decision structure with robustness against worst-case uncertainty $\min_{x\in X}c^{\top}x+\max_{\xi\in\mathcal{U}}\min_{y\in Y(x,\xi)}f(x,y,\xi)$ L-shaped method[104], nested C&CG[46][22], C&CG[106]
Tri-level stochastic robust optimization Combines stochastic and robust elements: outer stage handles here-and-now, middle stage captures uncertainty distribution/worst-case, inner stage executes corrective recourse $\min_{x}c^{\top}x+\mathbb{E}_{\xi}[\max_{u\in\mathcal{U}(\xi)}\min_{y}f(x,y,u)]$ Nested C&CG with progressive hedging[47]
Distributionally robust optimization Decisions are robust against the worst-case probability distribution within an ambiguity set, rather than a single known distribution $\min_{x\in X}c^{\top}x+\sup_{P\in\mathcal{P}}\mathbb{E}_{\xi\sim P}[f(x,\xi)]$, where $f(x,\xi)=\min_{y\in Y(x,\xi)}\{q^{\top}y:Wy+Tx\geq h\}$ C&CG with strong cutting planes[112]
Bi-level optimization Leader–follower structure where the upper level (leader) decision depends on the optimal response of the lower level (follower) $\min_{x\in X}c^{\top}x+F(x,y)$ s.t. $y\in\arg\min_{y\in Y(x)}f(x,y)$ Reformulation via KKT conditions[107], evolutionary algorithms[113]
VI-B Multi-type Hydrogen Facilities
To maximize the potential of utilizing hydrogen-related facilities against extreme events, multi-type hydrogen facilities could be considered. For example, Ameli et al.[108] proposed a resilient planning method to decide the capacity and location of HESS and hydrogen production facilities for integrated gas and electricity systems with HESSs considering deterministic scenarios under extreme weather events. Different from the above studies, Shao et al.[99] proposed a planning method to decide the capacity, location, and quantity of HESS, CCHP, and SHS in hydrogen-based microgrids with the consideration of uncertainties related to power supply, load demand, and device failure caused by disastrous events. The framework of the proposed planning method is shown in Fig. 10, where data sources related to renewable energy sources, load demand, and device failure are first prepared. Then, scenarios used for planning are generated, which consist of normal scenarios and emergency scenarios. Next, a two-stage risk-constrained stochastic optimization framework is adopted for problem formulation, where the investment decision stage belongs to the first stage, and the operation stage belongs to the second stage. After the investment plan is determined, it will be transmitted to the second stage. According to the generated scenarios, the corresponding normal-status operation or on-emergency operation is carried out, and the resulting performance metrics are used for the adjustment of the investment plan of the first stage. Note that risk constraints are considered to reduce the conservatism of the investment plan. Simulation results indicated that the proposed planning method can implement flexible tradeoff between economics and resilience. In [109], Xie et al. developed a planning problem model related to hydrogen-based seaport MESs, where normal operation scenarios and contingency operation scenarios are simultaneously considered. The model is formulated as a two-stage stochastic programming, where the first-stage decision is to decide the quantity and location of HRSs, fuel cell electric trucks, renewable generators, and remote control switches, and the second-stage decision is to decide scheduling power of each equipment under the given scenario. Simulation results indicated that the investment of resilience and low-carbon energy infrastructure can bring three aspects of benefits, i.e., power outage risk, system cost, and environmental conservation. In [110], Wang et al. proposed a risk-averse two-stage stochastic programming framework for the economic–resilient planning of hydrogen-enriched distribution networks with MHERs, where the first stage minimizes investment cost by co-siting and sizing stationary distributed energy resources and MHERs. To hedge long-run operational risk, a conditional-value-at-risk-based objective is introduced, and on-emergency corrective actions (including energy storage dispatch, dynamic re-routing of MHERs, and distribution feeder reconfiguration) are modeled to enhance system resilience in the second stage. Case studies demonstrated that proper configuration of stationary and mobile distributed energy resources reduces load-curtailment risk by more than 75% under comparable expenditure levels.
In order to deal with the worst-case cost under extreme events, Dong et al.[22] proposed a co-planning method for hydrogen-based microgrids and fuel-cell bus operation centers to decide the capacity, location, and quantity of HRS and FC electric bus considering stochastic line outages incurred by extreme weather events. Note that the proposed planning method is based on a nested C&CG algorithm, which can solve the formulated $\max-\min-\max$ tr-level optimization problem efficiently. In [111], Liu et al. investigated a hydrogen pipeline hardening and HESS allocation problem for electric-hydrogen network against extreme weather. Then, the above problem was formulated based on two-stage distributionally robust optimization framework, where the first stage focuses on pre-disaster decision under the given budget and the allowed hydrogen leakage risk, the second stage focuses on minimizing load shedding cost after events happen. The formulated problem was transformed into a two-stage MISOCP problem and was solved by C&CG problem. Simulation results indicated that the proposed strategy can minimize load shedding and control hydrogen pipeline leakage risk. Similar works could be observed in [112]. In [113], Wang et al. developed a resilient planning method for hybrid hydrogen and battery energy storage systems under extreme events based on particle swarm optimization and MILP, where particle swarm optimization is adopted to decide the system configuration (including installed location of fuel cell and battery; the rated power of fuel cell, battery, electrolyzer, and hydrogen tank; the capacity of battery) and the MILP related to the system operation cost minimization is solved by commercial solvers. In this work, parameter uncertainty and event uncertainty are not considered, while contingency modeling related to renewable energy reduction, load increase, and line capacity reduction is considered. Simulation results indicated that the hybrid energy system can improve the resilience index (i.e., load loss ratio) by 23.8% and 0.7%, compared with single battery energy storage or hydrogen energy storage, respectively.
Figure 10: The framework of the two-stage planning method
Summary: To show the differences of existing works more clearly, their main features are summarized in Table VI and Table VII. It can be observed that not all works considered three kinds of uncertainties simultaneously, e.g., extreme event uncertainty, event-incurred contingency uncertainty, and system parameter uncertainty. Moreover, when more complex HMESs are considered (e.g., electricity-hydrogen-gas systems), the number of decisions related to different components becomes larger. Consequently, the formulated problems have more complex forms, e.g., MISOCP and MINLP. In addition, there are six kinds of modeling frameworks for planning enhancement problem of HMESs, i.e., single-level deterministic programming, bi-level programming, two-stage stochastic programming, three-level robust programming, three-level stochastic robust programming, and three-level distributionally robust optimization. Their modeling explanations, mathematical formulation, and solution method are summarized in Table 8. In row 1, $x$ denotes all decision variables, $c^{\top}x$ denotes the investment cost, $f(x)$ denotes other cost except investment cost; In row 2, $y$ denotes decision variables of lower-level optimization problem with objective function $f(x,y)$, $F(x,y)$ denotes the operation cost in the upper-level optimization problem; In row 3, $Q(x,\xi)$ denotes the operation cost under the scenario $\xi$; In row 4, $f(x,y,\xi)$ denotes the operation cost under the scenario $\xi$; In row 5, $f(x,y,u)$ denotes the operation cost under a decision $u$, which belongs to $\mathcal{U}(\xi)$; In row 6, $p$ denotes a probability distribution, $f(x,\xi)$ denotes the optimal operation cost under the scenario $\xi$ and decision $x$. Finally, the above works mainly focus on simultaneous line outages and/or device faults incurred by natural extreme events, while cascading faults and cyber attacks are seldom considered[114].
VII Hydrogen-enabled Operation Enhancement for MESs
As mentioned above, three types of HMES operation response stages under extreme events are involved. According to the measures adopted in the above-mentioned three stages, existing works can be divided into two types, i.e., single-stage operation strategy and multi-stage operation strategy. The former means that the designed strategy is involved in one operational stage, while the latter means that the designed strategy is involved in multiple operational stages. In the following parts, the related works are introduced in detail.
VII-A Single-stage operation strategy
Many single-stage operation strategies have been developed in existing works. For example, Sharifpour et al.[85] proposed a preventive response strategy for a networked microgrid by integrating HESSs and demand response programs, which encourages replenishing HESSs and restricting non-essential loads before the impending main grid power outages incurred by extreme events. Similarly, Shahbazbegian et al.[84] proposed a preventive response approach to optimize operation cost and resilience metrics (e.g., electricity purchasing quantity, hydrogen tank level, and power loss) related to a microgrid with power-to-hydrogen systems when line outages caused by extreme events are considered. Although some advances have been made, the above methods assume that the future event attributes (e.g., power outage duration and starting time) are known. Moreover, parameter uncertainties related to load and renewable generation were neglected. To overcome such challenges, many preventive response strategies are designed with the consideration of uncertainty modeling[78][29]. For example, in [78], Yuan et al. proposed a risk-constrained day-ahead preventive response method for photovoltaic-penetrated power distribution networks equipped with power-to-hydrogen systems based on two-stage stochastic programming considering uncertainties in wind speed, solar radiation, active and reactive loads. Liu et al.[29] proposed a resilient day-ahead scheduling method for power distribution networks in the presence of uncertainties in renewable output, active/reactive demands, and disaster duration, which intends to prepare HRSs for electricity generation during power outage emergencies. In [30], Zhao et al. proposed a robust hydrogen-penetrated formation strategy against emergent power outages caused by extreme events with the help of HRSs and hydrogen transit while taking uncertainties in load and hydrogen demands, renewable generators, and temporal distances between HRSs into consideration.
[…]
Figure 11: The key idea of the proposed scheduling method for the hydrogen-based multi-energy supply microgrid
Figure 12: The illustration of an energy hub under threats of false data injection attacking
After the extreme events, some measures related to system topologies or resources could be adopted to restore load. For example, Zhu et al.[79] developed a resilience-oriented service restoration strategy for a power-hydrogen distribution system by taking some measures, e.g., network topology reconfiguration, scheduling of dispatchable generators and HESSs, and load prioritization. Similarly, Afsari et al.[74] proposed a service restoration strategy for microgrids with hydrogen storage based on network topology reconfiguration and the coordination of HESSs and hydrogen-based micro-gas turbines. Before network reconfiguration is performed, an HMES black-start is required. In [57], Lu et al. developed a black start strategy for hydrogen-integrated renewable grids with energy storage systems, where hydrogen fuel cells are used as black start resources. In addition to stationary HERs, MHERs could also be scheduled for restoration response. For example, Su et al.[80] proposed a restoration strategy for hydrogen-accommodated microgrids based on a two-stage stochastic programming framework, where the location allocation of mobile wind turbines is decided in the first stage according to the shortest-path information in the transportation network considering uncertainties in wind generation. Then, the expected costs of power outages will be minimized by scheduling mobile wind turbines and HESSs in microgrids jointly. In [76], Zhao et al. proposed a two-level restoration strategy for a distribution network with HRSs. As shown in Fig. 13, the distribution system operator schedules mobile energy sources (i.e., HTTs and HFCVs) to minimize the power imbalance among microgrids and reduce the unserved load in the upper level. Then, based on the allocation results, a microgrid energy management strategy is designed to minimize the system operation cost of each microgrid. Simulation results showed that the coordination of mobile energy sources and HFCVs can reduce the power imbalance and improve the self-healing ability of hydrogen-electricity systems. Different from above works, Wang et al.[120] proposed a restoration strategy for an integrated power and hydrogen distribution system with the consideration of power line and hydrogen pipeline outages caused by extreme events. When power outages happen, transmission line and hydrogen pipeline reconfiguration was first adopted to avoid islanding. Then, repair crews were dispatched to repair the fault lines and hydrogen pipelines. In addition, mobile battery-carried vehicles were dispatched to cooperate with repair crews to restore critical loads.
Figure 13: The framework of the two-level restoration strategy for a distribution network with HRSs
VII-B Multi-stage operation strategy
Although the above single-stage operation strategies can enhance HMES resilience, they cannot implement the coordination between different stages. To this end, some multi-stage operation strategies for HMESs have been proposed. For example, Haggi et al. proposed a proactive rolling horizon optimization-based method for hydrogen systems to enhance the power system resilience[50], which can implement the seamless switching between normal and emergency modes. Under normal mode, hydrogen systems assist the power grid by acting as a load. Once the extreme event time is known based on the forecast, the hydrogen tank is filled to make preparation for emergency mode. In [86], Liu et al. developed a risk-averse receding horizon optimization method to support the proactive and emergency operation of the industrial park based on integrated hydrogen-electricity-heat microgrids, where the measure of proactive operation includes increasing the reserve capacity of FCs and that of emergency operation is to reschedule components for the electricity and heat load survivability. In [82], Wang et al. proposed a rolling dispatch method to enhance the resilience of hydrogen-penetrated Antarctic energy systems under extreme weather. The framework of the rolling dispatch method can be observed in Fig. 14, where three stages could be identified, i.e., normal scheduling, alert scheduling, and emergency scheduling. Moreover, the above-mentioned stages will transfer from one to another according to different event states. To be specific, when the extreme event information is known based on forecasting techniques at slot $t+s$, the normal scheduling stage with horizon length $h$ will transfer to the alert scheduling stage with rolling horizon length $h+\tau$. Here, $\tau$ is the extreme event duration. When the event beginning is detected, the alert scheduling stage will transfer to the emergency scheduling stage with a rolling horizon length $\tau$. Once the event ending is detected, the alert scheduling stage will transfer to the normal scheduling stage. Since the function of alert scheduling is similar to that of preventive response (i.e., making preparation for emergency response), their roles are regarded as equivalent in Table VI. Similar rolling horizon optimization-based methods considering economic operation, proactive preparation, and emergency operation can be observed in [89].
Figure 14: The framework of the rolling dispatch method with an alert mechanism
TABLE IX: Summary of Operation Enhancement for HMES (Single-stage)
Ref. Events Uncertainties Demand types Operation stages Problem types and solving methods Measure types
[85] Main grid power outages caused by extreme events $\times$ Electricity Preventive MILP; CPLEX Type-A
[84] Main outages caused by extreme events $\times$ Electricity Preventive MINLP; Generalized Benders Decomposition Type-A
[94] Line outages and generator disruption $\times$ Electricity, hydrogen Preventive MINLP Type-A
[78] Main grid power outages caused by extreme events Wind speed, solar radiation, active and reactive loads Electricity, hydrogen Preventive Stochastic MILP, fuzzy method Type-A
[29] Main grid power outages caused by natural disaster Active and reactive demands and renewable generation Electricity, hydrogen Preventive Stochastic MILP; CPLEX Type-A
[30] Main grid power outages caused by extreme events Electricity and hydrogen demands, renewable generators, temporal distances between HRSs Electricity, hydrogen Preventive Stochastic MILP; Gurobi Type-B
[51] Main grid power outages Renewable energy and electricity prices Electricity, hydrogen Emergency Stochastic MILP, CPLEX Type-D
[121] Main grid power outages caused by extreme events $\times$ Electricity, hydrogen Emergency Fuzzy multi-objective MISOCP; Gurobi Type-D
[117] Line outages and generator trips caused by ice disasters Line outage probability Electricity, hydrogen Emergency MILP, CPLEX Type-E
[31] Main grid power outage and insufficient natural gas supply caused by extreme events $\times$ Electricity, gas Emergency MILP; Alternating direction method of multipliers Type-E
[18] Component failures caused by hurricanes $\times$ Electricity, gas, heat Emergency MILP, Frank-Wolfe algorithm Type-F
[115] Component outage caused by extreme events Loads, renewables, electric vehicle charging behaviors Electricity, heat, cooling Emergency Stochastic MILP, GAMS solver Type-H
[119] Extreme PV and wind turbine generation PV and wind turbine generation Electricity, heat, hydrogen Emergency Stochastic MILP, Gurobi/CPLEX Type-E, Type-G
[59] False data injection cyber-attacks on electricity and hydrogen demand $\times$ Electricity, heat, hydrogen Emergency MILP; CPLEX Type-G
[79] Distribution line and pipeline outages caused by extreme events $\times$ Electricity, heat, hydrogen Restoration MISOCP; Relaxation inducement Type-K
[74] Line outages caused by natural disasters Load demand Electricity Restoration Chance-constrained adaptive distributionally robust MILP; CPLEX Type-K
[57] Large-scale blackouts caused by extreme events $\times$ Electricity Restoration MILP Type-L
[80] Line outages caused by extreme weather events Wind forecast Electricity Restoration Stochastic MILP; Gurobi Type-I
[76] Line outages caused by extreme events Renewable outputs, electric and thermal loads Electricity, heat, hydrogen Restoration Stochastic MILP; Gurobi Type-J
[120] Power line and hydrogen pipeline outages caused by extreme events $\times$ Electricity, hydrogen Restoration MILP, Gurobi Type-M
TABLE X: Summary of Operation Enhancement for HMES (Multi-stage)
Ref. Events Uncertainties Demand types Operation stages Problem types and solving methods Hydrogen-related measure types
[86] Critical component contingencies caused by extreme events Network contingencies Electricity, heat Preventive, emergency Stochastic MISOCP; Receding horizon method Type-A, Type-D
[82] Outage of renewable energy units by storms Renewable energy Electricity, heat Preventive, emergency MILP; Gurobi Type-A, Type-D
[89] Main grid power outages and wind power plant contingencies caused by hurricanes Extreme event starting time Electricity, heat, hydrogen Preventive, emergency Rolling-horizon MILP; Gurobi Type-A, Type-D
[72] Line outages caused by hurricane Failed transmission lines, their failure, and repair times Electricity, hydrogen Preventive, emergency Stochastic multi-objective MILP Type-D
[23] Line outages and reduced PV generation caused by freezing rain diaster Line outages Electricity, heating, cooling, hydrogen Preventive, emergency MILP; Gurobi Type-A, Type-D
[11] Line outages caused by extreme weather Line faults Electricity Preventive, emergency Two stage robust optimization; C&CG Type-B, Type-F
[73] Line outages caused by hurricane Damaged transmission lines, hurricane time, and repair time Electricity, heat, hydrogen Preventive, emergency Stochastic MILP Type-C, Type-D
[77] Main grid disconnection and line outages caused by extreme events Line faults Electricity Preventive, emergency Tri-level MILP; Nested C&CG Type-C, Type-F
[81] Main grid power outages caused by extreme events Load, renewable energy, and travelling time Electricity, heat Preventive, restoration Stochastic MILP; Gurobi Type-D
[87] Lines/equipment failures incurred by hurricanes Lines/component failure probability Electricity Preventive, restoration Stochastic multi-objective programming; NSGA-II Type-C, Type-J
[75] Line outages caused by typhoons $\times$ Electricity, hydrogen Preventive, restoration MISOCP; CPLEX Type-A, Type-J, Type-K
[122] Main grid disconnection during emergencies $\times$ Electricity, heat, cooling, hydrogen Emergency, restoration MILP; CPLEX Type-G, Type-J
[21] Line outage caused by extreme events Renewable generation, load demands, and charging level of EVs Electricity, heat, cooling, hydrogen Emergency, restoration Bi-level MILP; Adaptive ADMM algorithm Type-G, Type-J
[123] Line outages and insufficient gas supply caused by extreme events Demand and renewable generation Electricity, heat, gas, hydrogen Emergency, restoration Stochastic MIQCP; Augmented ADMM algorithm Type-F, Type-D, Type-J
TABLE XI: Representative Hydrogen-related Operation Measures for HMES resilience enhancement
Stages Measure types Hydrogen-related measures Examples of specific measures
Preventive Type-A HESS pre-charging Increasing HESS level for reserve[50][75][82][84, 85, 86][23]
Type-B HTT scheduling HRS-based microgrid formation assisted by transport-based measures (e.g., HTTs and mobile hydrogen trucks) [30], pre-disaster HTT scheduling[11]
Type-C MHER location allocation Location allocation for HTTs[77], sizing and locating for HFCVs[87]
Emergency Type-D HESS discharging Hydrogen tank-driven fuel cell for independent power supply during outages[51], discharging batteries and HESS[89], HESS and CHP work at maximum power[82], discharging HESS for serving HFCVs[23], discharging HESS for power supply[86]
Type-E Natural gas pipeline-assisted hydrogen use, hydrogen pipeline scheduling Utilizing hydrogen enriched natural gas for fuel support[31], using pipelines as large-scale energy storage[119], using hydrogen stored in the pipeline for power support[117]
Type-F HTT/hydrogen truck scheduling Using mobile hydrogen trucks for critical load support[18][19], fuel cell power supply supported by hydrogen from HTTs[11], dynamic re-routing of HTTs[77]
Type-G Hydrogen demand adjustment Shedding non-critical hydrogen demand[119], restoring the integrity of compromised hydrogen demand data [59], HRSs adjust the charging schedule of HFCVs[21][122]
Type-H Hydrogen transformation Using hydrogen generated from renewables for synthetic natural gas production to satisfy demand[115]
Restoration Type-I Mobile hydrogen generation and storage Combined operation of mobile wind turbines and HESSs[80]
Type-J HFCV or hydrogen truck scheduling HFCV routing and dispatch[75][76][87], HFCV for power support[75], hydrogen fuel cell trucks are scheduled to HRSs for charging and dispatched to connection points for power supply[21][122]
Type-K HESS scheduling Scheduling HESS with other components[81], HESS discharging[74][75]
Type-L Hydrogen fuel cell scheduling Hydrogen fuel cells as black-start resources[57]
Type-M Hydrogen pipeline scheduling Hydrogen pipeline reconfiguration and repair[120]
Different from the above works, some studies considered the use of mobile energy sources for multi-stage operation. For example, Cai et al.[72] proposed a preventive response strategy to decide the location of mobile batteries in hydrogen fuel station-integrated power systems considering the uncertainties of failed transmission lines, their failure, and repair times. In [23], Chen et al. proposed a two-stage resilience enhancement method against the freezing rain disaster for the highway transportation energy system. In the first stage, preventive response is conducted to pre-charge storage systems (including heating ESS, cooling ESS, electric ESS, and HESS) so that stored energy can be discharged to satisfy multiple demands during both emergency and response stages. The optimization objective in this stage is to minimize economic operation costs and maximize the stored energy in all kinds of ESSs. In the second stage, emergency response is adopted to minimize multi-energy load shedding amount and operation cost. In [11], Yang et al. proposed a two-stage robust operation strategy to decide the location and hydrogen weight of HTTs considering the uncertainties in line outages. Once the uncertain contingency is discovered, HTTs and HRSs will be scheduled to minimize load shedding of electricity and hydrogen demand. In [73], Jordehi et al. considered using mobile energy sources and demand response programs to improve the resilience of the power system under extreme events. To be specific, the location of mobile energy sources was first decided considering the uncertainties in damaged lines, hurricane time, and repair time. Then, the joint scheduling of mobile energy sources, demand response programs, and energy hubs was conducted. In [77], Cao et al. investigated a resilient scheduling problem for electricity-hydrogen distribution networks considering line fault uncertainties. Then, a tri-level optimization-based scheduling strategy was proposed. The key idea of the proposed operation strategy is shown in Fig. 15, where two stages could be identified. In the preventive response stage, the locations of MHERs at different electricity-hydrogen nodes are decided. Then, the rerouting of MHERs with the help of the transportation network and network reconfiguration under the worst-case damage scenario are optimized jointly in the emergency response stage. Given guaranteed load survivability, the proposed scheduling strategy effectively reduces unit reinforcement cost. In addition to the combination of preventive response and emergency response, some studies considered the coordination of preventive response and restoration response. For example, [81] et al. investigated a resilience enhancement problem of an electricity-transport coupled system considering power outages caused by extreme events under the uncertainties in load, renewable energy, and electric vehicle commuting time. In the preventive stage, a pre-location of electric vehicle was conducted. In the restoration stage, the scheduling of electric vehicles and HESS was implemented. Simulation results showed that the proposed approach offers benefits in aspects of resilience enhancement and supplying critical loads. Similarly, Tang et al.[87] proposed a sizing and location allocation strategy for electric emergency power supply vehicles and hydrogen fuel cell emergency power supply vehicles so as to achieve a balance between resilience and economics. In [75], Xie et al. proposed a method to enhance power distribution network resilience based on HESSs and HFCVs. In the preventive stage, pre-deployment of HESSs and HFCVs within the electric-road network system was conducted. In the restoration stage, HESSs and HFCVs are scheduled to minimize load power losses and economic costs. Simulation results indicated that network resilience and recovery performance can be enhanced by adopting hydrogen technologies. In order to enhance the resilience of system resilience after natural disasters, Li et al. developed a two-layer economic resilience model for distribution network. Specifically, mobile battery storage units are considered and HRSs adjusted the charging schedule of HFCVs during emergency response. In the restoration response, hydrogen fuel cell trucks were dispatched to connection points for electricity supply. Similar measures could be found in [21]. In [123], Chen et al. proposed a self-healing approach for coupled electricity-gas-heat-hydrogen networks, which consists of two parts, i.e., hydrogen fuel cell trucks with storage were deployed to increase hydrogen energy storage level under emergency conditions, and such trucks were dispatched to connection points for electricity supply in restoration response. Simulation results indicated that the proposed approach can harness maximum capacities from smart prosumers for reducing load shedding and improving the infrastructure resilience.
Figure 15: The framework of resilient scheduling for electricity-hydrogen distribution networks
Summary: For easy understanding of existing works related to resilient operation for HMESs, we summarize their characteristics in Table IX and Table X, and list hydrogen-related measures in each response stage in Table XI. It can be observed that hydrogen plays an important role in every response stage, such as pre-charging HESS in the preventive response stage, hydrogen for fuel support via natural gas pipeline in the emergency response stage, and scheduling MHERs in the restoration response stage. Moreover, existing works mainly focus on single-stage or two-stage operation methods. Few of them consider the coordination operation among three kinds of response stages. In addition, hydrogen pipeline-related operational measures in emergency response and restoration response could be observed. The reason is that hydrogen pipelines can serve as large-scale energy storage solutions and play a key role in resilience enhancement. As mentioned in existing work[119],[124], the pressure of a hydrogen pipeline can be adjusted dynamically by changing the quantity of hydrogen within it. By exploiting such elasticity, many resilience enhancement measures related to hydrogen pipelines could be adopted. For example, MESs can temporarily store hydrogen in the pipeline when there are extra renewable energy but limited hydrogen tank space. When extreme event happens, the hydrogen pipeline can release hydrogen to drive fuel cell for energy supply. Such linepack utilization is helpful for enhancing resilience of MESs[125]. Moreover, during extreme events (such as main grid outages or natural gas supply interruptions), hydrogen pipelines can be directionally dispatched to recover hydrogen supply to critical loads such as hospitals and communication base stations. Finally, most contingencies in hydrogen-enabled resilience enhancement for MESs in existing works are related to main grid power outages, component outages, renewable generation reduction, and load surge, cyber-attack related contingencies should be paid more attention, which are also very common in cyber-physical energy systems[126, 127, 128, 129].
VIII Research Gaps and Future Directions
Although some advances have been made in existing works, several important research gaps are identified and possible directions are provided below.
VIII-A Comprehensive Design of HMES Resilience Metrics
Current resilience metrics in HMESs are often borrowed from power systems and MESs, such as LSR, ELNS, ERL, and WSLL. These metrics are primarily designed to quantify the extent of unmet energy demand in electricity, heating, and cooling during and after extreme events. However, these metrics fail to capture physical characteristics and functional complexities of hydrogen systems, such as hydrogen availability and sector-coupled interactions. To be specific, hydrogen availability is not binary and it is strongly influenced by pressure levels within pipelines and storage units. For example, fuel cells require hydrogen at a minimum operating pressure (e.g., >>1.5 bar). Thus, insufficient pressure, even with sufficient quantity of hydrogen present, leads to functional failure. Consequently, such “hidden” functional failure may result in serious social impacts under some extreme events, e.g., the failed operation of life-support equipment in an emergency hospital supported by hydrogen fuel cells may result in the loss of many patients’ lives. Additionally, HMESs rely on sector-coupled interactions (e.g., excess electricity is used to generate hydrogen via electrolysis, which is later used for electricity generation or heating). A failure in the hydrogen link breaks this coupling, degrading system functionality without necessarily causing measurable load loss. Therefore, existing HMES resilience metrics should be extended to incorporate hydrogen-specific attributes. A possible solution is to develop comprehensive resilience metrics by considering hydrogen-specific attributes (e.g., hydrogen system availability ratio and pressure recovery time), load loss, the number of customers affected, and carbon resilience jointly.
VIII-B Temporally-correlated Event-targeted Scenario Generation for Planning
In existing works on planning methods for the resilience enhancement of HMESs, uncertainties related to system parameters (e.g., distributed generation output and versatile demands) are seldom considered, or just several representative days are considered. In fact, the contingency duration incurred by an extreme event (e.g., extreme snow and freezing ice disasters) may last for several days[10][131]. Thus, the representative day is not long enough to represent the contingency uncertainty. Moreover, when taking system parameters into consideration, their correlations should be utilized to generate more representative and practical scenarios. For example, in some places, when heat demand is needed in winter, the corresponding cold demand may be not required. Similarly, when cold demand is needed in summer, the corresponding heat demand may be negligible[132]. In addition, extreme events generally have spatial features and their uncertainties should also be considered for scenario generation[47]. In summary, how to generate temporally-correlated and event-targeted scenarios for planning method design requires further investigation. A possible solution for generating temporally-correlated scenarios is to generate normalized multi-parameter vector scenarios according to the duration of extreme events and condense them using scenario reduction methods. Finally, the condensed normalized multi-parameter vectors are recovered using denormalization operations.
VIII-C Multi-type Temporal-spatial Cyber-physical Contingency Modeling under Compound Extreme Events
Most existing event-oriented contingency models focus on static, single-type faults, e.g., main grid power outages, simultaneous line failures, and simultaneous device failures. However, due to the coupling features of MESs, multi-type faults are more common under extreme events, e.g., simultaneous electrical network outages, natural gas network outages, and renewable generator failures[133]. In addition, with the progress of extreme events, the component failure probabilities may also change, e.g., with the increase of ice thickness, component failure probabilities under different time intervals are different[23]. Consequently, sequential multi-type failures may happen. Similarly, spatial dynamics of contingencies incurred by a hurricane could be observed in practice[18]. Furthermore, when compound extreme events happen (e.g., typhoons, earthquakes, freezing ice disasters, tropical-cyclone-blackout-heatwave hazards, and cyber attacks)[7], the resulting temporal and spatial dynamics of faults are more complex[134]. Above all, hydrogen-specific cyber vulnerabilities (e.g., false data injection attacks) are seldom considered. Thus, multi-type temporal-spatial cyber-physical contingency modeling under compound extreme events needs further investigation. A possible solution for such modeling is to use spatially-aware dynamic Bayesian networks[135][136] and Petri-net[137] since the former can capture the probabilistic dependencies of spatially-correlated component states under extreme events and the latter can capture the MES operation logic related to energy state and energy transition.
VIII-D Multi-network Multi-timescale Coordination Planning and Operation
Most existing works mainly focus on enhancing the resilience of HMESs related to electricity demand and hydrogen demand. However, two aspects are neglected. Firstly, multi-timescale feature existing in electric-hydrogen systems is not considered. To be specific, slow dynamics of hydrogen systems and the fast response of electric systems are coupled with each other. For instance, hydrogen pipelines exhibit response delays and they can not respond instantly to changes in flow or pressure, i.e., the pipeline system does not reach a new equilibrium immediately after a disturbance, and it takes time for changes in input or demand to affect the system’s operation. Such delay may span from minutes to hours. In addition to slow dynamics, hydrogen pipeline can be used to store hydrogen temporarily due to the compressibility of hydrogen, which can help to absorb fluctuations in renewable energy generation or load demand without the immediate need for additional storage tanks. Compared with hydrogen tank/pipelines, large-scale hydrogen energy storage systems (e.g., salt-cavern hydrogen storage) have slower dynamics, which are used as SHSs to adjust the weekly/monthly source-load imbalance[54]. In contrast, to capture the minute-level fluctuations of renewable energy and load demand, electricity system could be scheduled at minute timescale. Thus, multi-timescale modeling should be considered in the planning and operation of hydrogen-enabled MESs. Secondly, the same attention should be paid to providing heat demand and gas demand with high service quality under extreme events, since such demands may exceed 50% of total energy demand in a country[108], e.g., almost half of the energy demand in the U.K. is attributed to heat demand. When HMESs with versatile energy demands are considered, more strong couplings of power/gas/heat networks would be incurred. To manage such systems of systems efficiently, multi-network coordination is required. since enhancing the HMES resilience requires the use of many MHERs (e.g., HTTs, HFCVs, and HFCBs) and other mobile electric energy resources (e.g., mobile wind turbines and electric vehicles), four-network (i.e., power/gas/heat/transportation networks) coordination in planning and operation deserves further investigation. To support multi-timescale coordination operation, a possible solution is to utilize hierarchical MPC or hierarchical deep reinforcement learning for multi-timescale coordination operation, e.g., deciding the quantity of MHERs at day timescale, scheduling MHERs in emergency response at the hourly timescale, and scheduling electric equipment at the 5-minute timescale.
VIII-E Low-carbon and Resilient Planning and Operation
As highlighted in [138, 139, 140], achieving both low-carbon development and resilience is essential for the future of energy networks, including HMESs discussed in this paper. However, most existing studies on HMES resilience enhancement have largely overlooked the low-carbon objective. For example, in [28], carbon emission costs were included in the objective function of a planning model for a hydrogen-integrated airport energy system, helping to reduce system-level emissions. Despite such efforts, focusing solely on low-carbon strategies is insufficient, as the ultimate target for future energy systems is to achieve net-zero carbon emissions [141, 142]. Therefore, future planning and operational strategies for HMESs must jointly address both net-zero carbon targets and system resilience. To tackle this dual-objective challenge under uncertainty, advanced artificial intelligence methods can be employed. One promising solution is multi-objective deep reinforcement learning [143, 144, 145, 146], which enables coordinated decision-making across multiple objectives, timescales, and stages. Another solution is to use deep reinforcement learning-assisted evolutionary algorithms [147, 148], where deep reinforcement learning helps dynamically control the crossover probability between individuals at different levels and the mutation intensity for individuals, thereby improving the quality of evolved offspring and enhancing global search performance.
VIII-F LLM-assisted Whole-Process Resilience Enhancement for HMESs
LLMs (e.g., GPT-4) have demonstrated remarkable capabilities in understanding, synthesizing, and reasoning across vast domains of unstructured and structured information[149, 150]. Their contextual awareness, semantic reasoning, and adaptability position them as promising tools for supporting the whole-process resilience enhancement of HMESs. Firstly, the current resilience metrics failed to capture the hydrogen-related attributes, e.g., hydrogen availability, tank/pipeline pressure recovery, and sector-coupled response. In contrast, LLMs can extract, analyze, and generate multi-dimensional, cross-sectoral resilience indicators by mining technical literature, policy documents, and incident reports, providing a foundation for standardizing hydrogen-specific evaluation frameworks. Secondly, contingency modeling under extreme events in HMESs is hindered by the complexity of hydrogen-related subsystems, while LLMs can contribute to contingency modeling by constructing structured fault trees and scenario libraries based on natural language descriptions, enabling the automatic generation of probabilistic or rule-based failure scenarios under diverse stress conditions. Thirdly, from a planning enhancement perspective, LLMs can support scenario generation and strategic siting of hydrogen assets by reasoning over multi-objective (e.g., operation cost, carbon emission, and resilience) trade-offs. Finally, from the operation enhancement perspective, LLMs offer potential in assisting the dynamic dispatch and routing of MHERs since they can interpret complex, multi-modal inputs (ranging from sensor data to emergency protocols) and contribute to human-in-loop decision.
IX Conclusions
This article presents the first comprehensive review of hydrogen-enabled resilience enhancement for MESs from the perspective of planning and operation. Firstly, advantages and challenges of adopting hydrogen to enhance resilience of MESs are introduced. Then, it provides a comprehensive resilience enhancement framework for HMESs. Under the proposed framework, the widely used resilience metrics and event-oriented contingency models are summarized. Then, it classifies the hydrogen-enabled planning measures for resilience enhancement of HMESs according to the type of hydrogen-related facilities and summarizes the typical planning problem formulation frameworks. Next, it categorizes the hydrogen-enabled operation measures for the resilience enhancement of HMESs according to three types of operation response stages and summarizes the role of hydrogen in each response stage. Finally, several research gaps and future directions are identified, i.e., (1) Comprehensive HMES resilience metrics with the consideration of hydrogen-specific attributes (e.g., hydrogen system availability ratio and pressure recovery time), load loss, the number of customers affected, and carbon resilience should be designed; (2) Generating temporally-correlated event-targeted scenario generation for planning based on normalized multi-parameter vector scenarios and scenario reduction methods; (3) Achieving multi-type temporal-spatial contingency modeling under compound extreme events based on spatially-aware dynamic Bayesian networks and Petri-net; (4) Implementing multi-network multi-timescale coordination planning and operation based on hierarchical MPC or hierarchical deep reinforcement learning; (5) Designing advanced artificial intelligence-based multi-objective optimization methods for the planning and operation of low-carbon and resilient HMES; (6) Adopting LLM-assisted whole-process resilience enhancement for HMESs in aspects of proper resilience metrics, event-oriented contingency modeling, scenario generation and smart decision in planning process, and MHER scheduling and routing in operation process. While the above-mentioned technical strategies are essential for enhancing the resilience of HMESs, business and market models also play a crucial role. Representative models include hydrogen-as-a-service, priority hydrogen supply contracts, and shared hydrogen pipeline infrastructure. For instance, by leasing hydrogen fuel cell trucks under a hydrogen-as-a-service model, MESs can gain black-start capabilities and rapidly restore critical loads during unexpected power outages without requiring large upfront investments. Likewise, priority hydrogen contracts can ensure uninterrupted supply to critical infrastructure (e.g., hospitals or water treatment plants) during extreme events. Moreover, due to the high capital cost of hydrogen pipeline infrastructure, shared investment models among multiple MES operators can promote cost-effective deployment while improving cross-system resilience.

### v3

@@
 Abstract
-Ensuring resilience in multi-energy systems (MESs) becomes both more urgent and more challenging due to the rising occurrence and severity of extreme events (e.g., natural disasters, extreme weather, and cyber-physical attacks).
-Among many measures of strengthening MES resilience, the integration of hydrogen shows exceptional potential in cross-temporal flexibility, cross-spatial flexibility, cross-sector flexibility, and black start capability.
-Although many hydrogen-enabled MES resilience enhancement measures have been developed, the current literature lacks a systematic overview of hydrogen-enabled resilience enhancement in MESs.
-To fill the research gap, this paper provides a comprehensive overview of hydrogen-enabled MES resilience enhancement.
-First, advantages and challenges of adopting hydrogen in MES resilience enhancement are summarized.
-Then, we propose a resilience enhancement framework for hydrogen-enabled MESs.
-Under the proposed framework, existing resilience metrics and event-oriented contingency models are summarized and discussed.
-Furthermore, we classify hydrogen-enabled planning measures by the types of hydrogen-related facilities and provide some insights for planning problem formulation frameworks.
-Moreover, we categorize the hydrogen-enabled operation enhancement measures into three operation response stages: preventive, emergency, and restoration.
-Finally, we identify some research gaps and point out possible future directions in aspects of comprehensive resilience metric design, temporally-correlated event-targeted scenario generation, multi-type temporal-spatial cyber-physical contingency modeling under compound extreme events, multi-network multi-timescale coordinated planning and operation, low-carbon resilient planning and operation, and large language model-assisted whole-process resilience enhancement.
-Index Terms:
-Multi-energy systems, hydrogen-enabled resilience enhancement, resilience metrics, contingency modeling, planning, operation, extreme events, low-carbon
+Ensuring resilience in multi-energy systems (MESs) has become increasingly urgent and challenging due to the growing frequency and severity of extreme events, such as natural disasters, extreme weather, and cyber-physical attacks.
+Among the various approaches to enhancing MES resilience, hydrogen integration offers significant potential in cross-temporal, cross-spatial, and cross-sector flexibility, as well as black-start capability.
+Although considerable efforts have been devoted to this area, a systematic review of resilience enhancement in hydrogen-enabled MESs is still lacking.
+To address this gap, this paper presents a comprehensive review of hydrogen-enabled MES resilience enhancement.
+First, advantages, vulnerabilities, and challenges related to hydrogen-enabled MES resilience enhancement are summarized.
+Next, a resilience enhancement framework for hydrogen-enabled MESs is proposed, based on which existing resilience metrics and event-oriented contingency models are reviewed and discussed.
+Planning measures are then classified according to the types of hydrogen-related facilities, together with uncertainty handling methods, scenario generation methods, and planning problem formulation frameworks.
+In addition, operational enhancement measures are categorized into three response stages: prevention, emergency response, and restoration.
+Finally, research gaps are identified and future directions are discussed, including comprehensive resilience metric design, advanced extreme-event scenario generation, spatiotemporal cyber-physical contingency modeling under compound extreme events, coordinated planning and operation across multiple networks and timescales, low-carbon resilient planning and operation, and large language model-assisted whole-process resilience enhancement.
 I Introduction
@@
 To be specific, HERs refer broadly to the various technologies and systems that produce, store, transport, and utilize hydrogen as an energy carrier, including stationary HERs (e.g., electrolyzers, hydrogen tanks, hydrogen pipelines, hydrogen fuel cell-based conversion systems, and hydrogen refueling stations (HRSs)) and mobile HERs (MHERs), which include hydrogen tube trailers (HTTs)[11, 12, 13, 14, 15, 16, 17], mobile hydrogen trucks[18, 19, 20], mobile fuel cell trucks[21], hydrogen fuel cell buses (HFCBs)[22], hydrogen fuel cell vehicles (HFCVs)[23].
-The above HERs can offer distinguishing features for enhancing MES resilience in aspects of energy supply and carbon emission[24, 25, 26, 27], i.e., cross-temporal flexibility, cross-spatial flexibility, cross-sector flexibility, and black start capability.
+The above HERs can offer distinguishing features for enhancing MES resilience in aspects of energy supply and carbon emission as shown in Fig.
+1[24, 25, 26, 27], i.e., cross-temporal flexibility, cross-spatial flexibility, cross-sector flexibility, and black start capability.
 Despite the promising potential of hydrogen integration, developing effective resilience enhancement strategies for hydrogen-enabled MESs (HMESs) involves significant complexity across multiple dimensions.
@@
 Second, modeling extreme event-related HESS contingencies requires careful representation of multi-stage processes (such as electrolysis, compression, storage, and reconversion) and each process is subject to distinct cyber and physical operational risks.
-Third, designing the schemes of hydrogen-related infrastructure planning and HMES operation faces solving challenges caused by uncertainties, mixed integer variables, multi-timescale variables, temporal-spatial coupling constraints, nonlinear and nonconvex constraints, multi-energy coupling constraints, and multi-objective function.
+Third, designing the schemes of hydrogen-related infrastructure planning and HMES operation faces solving challenges caused by uncertainties, mixed integer variables, multi-timescale variables, temporal-spatial coupling constraints, nonlinear and nonconvex constraints, multi-energy coupling constraints, and multi-objective functions.
 To overcome the above challenges, researchers have defined many resilience metrics, developed some contingency models, and proposed many planning and operation methods for enhancing the resilience of HMESs[28, 29, 30, 31, 32].
 However, there is a lack of a systematic overview of hydrogen-enabled resilience enhancement for MESs.
-To fill the research gap, we provide a comprehensive overview on leveraging hydrogen to enhance the resilience for MESs.
+To fill the research gap, we provide a comprehensive overview of leveraging hydrogen to enhance the resilience of MESs.
 Although there have been many review works related to power grid resilience enhancement[34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 33] with different perspectives (e.g., networked microgrids, event timeline of enhancement, cyber-physical interdependence, complex network, and cross-domain multilayer architecture), they neglect to consider the HERs.
@@
 To the best of our knowledge, our work is the first survey dedicated to hydrogen-enabled resilience enhancement for MESs.
-The main contributions of this paper can be summarized below.
+Figure 1: Distinguishing features of hydrogen for enhancing MES resilience.
+Note that cross-temporal flexibility refers to the ability to store hydrogen over long durations and shift energy use from periods of surplus to periods of shortage; cross-spatial flexibility refers to the capability to transport hydrogen across regions to balance spatially uneven supply and demand; cross-sector flexibility refers to the interoperability among the electricity, heat, gas, and transport sectors through hydrogen conversion and blending; and black-start capability refers to the ability of an HMES system to self-start and restore electricity supply to critical infrastructure (e.g., hospitals, emergency shelters, and communication networks) without relying on external grid support during extreme events.
+The main contributions of this paper are summarized below.
   1. 1.
     We summarize the advantages and challenges of adopting hydrogen in MES resilience enhancement.
-Specifically, distinguishing features of hydrogen for resilience enhancement, resilience potential and security issues of hydrogen-related technologies, and optimization challenges of hydrogen-enabled planning and operation enhancement are summarized.
+Specifically, distinguishing features of hydrogen for resilience enhancement, resilience potential and security issues of hydrogen-related technologies, and optimization challenges of hydrogen-related planning and operation are summarized.
   2. 2.
@@
     We classify the planning enhancement of HMES resilience from two aspects, i.e., single-type hydrogen facility and multi-type hydrogen facility.
-The considered hydrogen-related facilities consist of HESSs, hydrogen pipelines, HRSs, HFCBs, fuel cell-driven combined cooling, heat, and power (CCHP), and hydrogen-related production facilities.
-Moreover, the insights in aspects of uncertainty representation and formulation framework for HMES planning are summarized.
+The hydrogen-related facilities considered include HESSs, hydrogen pipelines, HRSs, HFCBs, fuel cell-driven combined cooling, heat, and power (CCHP), and hydrogen-related production facilities.
+Moreover, we summarize the classifications of uncertainty handling methods, scenario generation methods, and problem formulation frameworks for HMES planning.
   4. 4.
@@
   5. 5.
-    We identify six research gaps and point out future directions in aspects of comprehensive resilience metric design, temporally-correlated event-targeted scenario generation, multi-type temporal-spatial cyber-physical contingency modeling under compound extreme events, multi-network multi-timescale coordinated planning and operation, low-carbon resilient planning and operation, and large language model-assisted whole-process resilience enhancement.
+    We identify six research gaps and point out future directions in aspects of comprehensive resilience metric design, advanced extreme-event scenario generation, multi-type temporal-spatial cyber-physical contingency modeling under compound extreme events, multi-network multi-timescale coordinated planning and operation, low-carbon resilient planning and operation, and large language model-assisted whole-process resilience enhancement.
 The rest of this paper is organized as follows.
@@
 ESS Energy storage system
+GAN Generative adversarial networks
 GSHP Ground-source heat pump
@@
 WSLL Weighted sum of load loss
-II Advantages and Challenges of Adopting Hydrogen in MES Resilience Enhancement
-In this section, we summarize the advantages and challenges of adopting hydrogen in MES resilience enhancement.
-To be specific, the distinguishing features of hydrogen for MES resilience enhancement are firstly provided.
-Then, from the perspective hydrogen supply chain, resilience potential and security issues related to hydrogen-related technologies are summarized.
-Finally, the optimization advantages in planning enhancement and operation enhancement of MES resilience are analyzed.
-Figure 1: Distinguishing features of hydrogen in MES resilience enhancement
+II Advantages, Vulnerabilities, and Challenges Related to Hydrogen-enabled MES Resilience Enhancement
+In this section, the distinguishing features of adopting hydrogen for MES resilience enhancement are firstly provided.
+Then, from the perspective of the hydrogen supply chain, potentials and vulnerabilities related to hydrogen technologies in resilience enhancement are summarized.
+Finally, optimization challenges associated with resilience-oriented planning and operation of hydrogen-enabled MESs are analyzed.
 II-A Distinguishing Features of Hydrogen for MES Resilience Enhancement
@@
 II-A1 Cross-temporal Flexibility
-Hydrogen can provide cross-temporal flexibility[46, 47, 48].
+This feature refers to the ability to store hydrogen over long durations and shift energy use from periods of surplus to periods of shortage[46, 47, 48].
 Unlike traditional battery energy storage systems, hydrogen can be stored at a much larger scale and over longer durations, e.g., underground salt-cavern hydrogen storage can be used to provide cross-seasonal low-cost and large-scale hydrogen storage[49].
@@
 In addition, when produced from renewable energy, hydrogen improves the consumption rate of renewable energy.
-By converting excess electricity into storable hydrogen using seasonal energy storages (SHSs) and conventional HESSs, MESs can buffer supply-demand mismatches more effectively and operate with greater carbon resilience (i.e., an energy system’s ability to maintain or recover to its current and predicted emission level during and after shocks such as renewable intermittency, public health emergencies, or climate-related disruptions)[53].
-As shown in [54], 100% self-sufficiency and zero carbon emission can be simultaneously achieved in distant oceanic islands with abundant renewable energy resources by coordinating SHSs and HESSs.
+By converting excess electricity into storable hydrogen using seasonal hydrogen storage (SHS) and conventional HESSs, MESs can buffer supply-demand mismatches more effectively and operate with greater carbon resilience (i.e., an energy system’s ability to maintain or recover to its current and predicted emission level during and after shocks such as renewable intermittency, public health emergencies, or climate-related disruptions)[53].
+As shown in [54], 100% self-sufficiency and zero carbon emissions can be simultaneously achieved in distant oceanic islands with abundant renewable energy resources by coordinating SHSs and HESSs.
 II-A2 Cross-spatial Flexibility
-Hydrogen can provide cross-spatial flexibility[46, 47, 48].
+This feature refers to the capability to transport hydrogen across regions to balance spatially uneven supply and demand[46, 47, 48].
 For example, hydrogen can be flexibly transported across regions through hydrogen pipelines, HTTs, or liquid hydrogen trucks, making it possible to allocate energy geographically according to regional demand and supply conditions.
@@
 Similarly, mobile hydrogen refueling trucks can be deployed to support remote or damaged areas after extreme weather events, such as floods, by providing temporary hydrogen fuel for microgrids and emergency vehicles.
-By bridging spatial gaps between supply and demand, hydrogen transportation infrastructure allows MESs to dynamically reallocate resources, maintain critical services, and prevent cascading failures in interconnected energy networks, resulting in an improved energy resilience.
+By bridging spatial gaps between supply and demand, hydrogen transportation infrastructure allows MESs to dynamically reallocate resources, maintain critical services, and prevent cascading failures in interconnected energy networks, resulting in improved energy resilience.
 II-A3 Cross-sector Flexibility
-Hydrogen plays a crucial role in enhancing the interoperability of energy systems by acting as a versatile bridge between multiple energy carriers, including electricity, heat, gas, and transport.
-For example, the power-to-hydrogen technologies enable surplus electricity from renewable sources like wind and solar to be converted into hydrogen, which can then be stored and later used for electricity and heat generation or as a fuel for HFCVs in transportation sector.
+This feature refers to the interoperability among the electricity, heat, gas, and transport sectors through hydrogen conversion and blending[31].
+For example, the power-to-hydrogen technologies enable surplus electricity from renewable sources like wind and solar to be converted into hydrogen, which can then be stored and later used for electricity and heat generation or as a fuel for HFCVs in the transportation sector.
 In addition, hydrogen can be converted to synthetic methane or blended with natural gas for transportation[31],[55],[56], which can reduce dependence on pure natural gas.
@@
 Based on dynamic reconfiguration, hydrogen can facilitate resource sharing across different energy vectors, especially during extreme events.
-In such cases, if one subsystem experiences a failure, hydrogen can provide an alternative energy pathway, thereby ensuring the continuous operation of essential services and improving the MES resilience.
+In such cases, if one subsystem experiences a failure, hydrogen can provide an alternative energy pathway, thereby ensuring the continuous operation of essential services and improving MES resilience.
 II-A4 Black Start Capability
-Hydrogen plays a critical role in safeguarding key infrastructures (e.g., hospitals, emergency shelters, and communication networks) during extreme events.
+This feature refers to the ability of an HMES system to self-start and restore electricity supply to critical infrastructure (e.g., hospitals, emergency shelters, and communication networks) without relying on external grid support during extreme events[57].
 In the event of prolonged power outages or natural gas supply failures, hydrogen-based infrastructure (such as proton exchange membrane fuel cells and hydrogen tanks) can operate off-grid and provide autonomous “black start” capabilities[57], significantly enhancing the system’s ability to self-recover without external intervention during extreme events.
-According to [57], black start should be implemented before network reconfiguration and load restoration could be done.
+According to [57], black start should be implemented before network reconfiguration and load restoration are performed.
 Compared with black-start diesel generators, hydrogen fuel cells produce zero local emissions, operate quietly, and offer higher reliability with fewer maintenance requirements.
@@
 3 Cross-sector flexibility Interoperability between electricity, heat, gas, and transport sectors via hydrogen conversion and blending Provides alternative energy pathways when one sector fails Facilitates green fuel substitution and decarbonization in multiple sectors
-4 Black start capability Off-grid operation of hydrogen fuel cells and tanks to restart or sustain critical infrastructure during outages Ensures rapid recovery and maintains essential services in blackout scenarios limited enhancement
+4 Black start capability Off-grid operation of hydrogen fuel cells and tanks to restart or sustain critical infrastructure during outages Ensures rapid recovery and maintains essential services in blackout scenarios Zero carbon emission compared with diesel generators
 II-B Resilience Potential of Hydrogen-related Technologies
-In the hydrogen supply chain, there are many hydrogen production pathways (e.g., electrolyzer, or steam methane reforming [58]), hydrogen storage manners (e.g., salt-cavern hydrogen storage, hydrogen tank), transportation methods (e.g., pipelines, trucks, or tube trailers), and hydrogen-to-power technologies (e.g., proton exchange membrane fuel cells, solid oxide fuel cells, and turbines), which offer different resilience capabilities in aspects of black start capability, off-grid operability, and response time, as summarized in Table III.
-By coordinating different hydrogen-related technologies, synergies can be achieved.
-For example, by combining proton exchange membrane fuel cell and on-site hydrogen tanks, fast, clean, and autonomous black start could be implemented for key infrastructures (e.g., hospitals).
-electrolyzers, salt cavern storage, pipelines, and hydrogen turbines, large-scale grid restoration could be implemented.
-Thus, a careful selection of hydrogen-related technologies tailored to the specific application scenarios is required for designing hydrogen-enabled MESs.
-TABLE III: Resilience Potential of Hydrogen-related Technologies
-Category Technology Black start capability Off-grid operability Response time Resilience Potential Summary
-Hydrogen production Electrolyzer No Yes (if coupled with renewables) Moderate (minutes) Enables distributed hydrogen production; useful in islanded scenarios but cannot self-start
-Hydrogen production Steam methane reforming No No Slow (hours) Suitable for centralized supply with high continuity, but lacks flexibility and emergency use
-Hydrogen storage Salt cavern storage No No Moderate (infrastructure-dependent) Suitable for seasonal, large-scale storage; requires energy infrastructure for operation
-Hydrogen storage High-pressure tank No Yes Fast (minutes) Modular and deployable; enables local backup supply and fast dispatch during disruptions
-Hydrogen transport Hydrogen pipeline No No Fast (seconds-minutes) Enables continuous hydrogen supply, supports regional resilience, but lacks autonomous function
-Hydrogen transport HTT No Yes Moderate (minutes-hours) Highly mobile for backup fueling in off-grid or emergency zones
-Hydrogen utilization Proton exchange membrane fuel cell Yes Yes Fast (seconds-minutes) Ideal for critical load support, mobile/emergency backup, and black start
-Hydrogen utilization Solid oxide fuel cell No (needs heat-up) Yes Slow (1-2 hours) Efficient and stable for long-term off-grid power but slow to start
-Hydrogen utilization Hydrogen turbines Yes (some models) Yes (with fuel supply) Moderate (minutes) Suitable for grid-scale backup and restoration, with moderate black start capability
-II-C Security Issues of Hydrogen-related Technologies
-Although resilience potential can be offered by hydrogen-related technologies in production, storage, transportation, and utilization, their security vulnerabilities also exist in each stage.
-During the production stage, especially in electrolysis and steam methane reforming, the risk of hydrogen leakage and explosion is significant due to high operating pressures and temperatures.
-During the storage stage, compressed hydrogen tanks face threats such as high-pressure rupture, while cryogenic liquid hydrogen presents risks of rapid venting.
+In the hydrogen supply chain, various hydrogen production pathways (e.g., electrolyzers and steam methane reforming [58]), storage options (e.g., salt-cavern hydrogen storage and hydrogen tanks), transportation modes (e.g., pipelines, trucks, and tube trailers), and hydrogen-to-power technologies (e.g., proton exchange membrane fuel cells, solid oxide fuel cells, and hydrogen turbines) provide different resilience capabilities in terms of cross-temporal flexibility, cross-spatial flexibility, cross-sector flexibility, and black-start capability, as summarized in Table III.
+The basis for the qualitative ratings (i.e., high, medium, and not obvious) is explained in the last column of the table.
+For example, salt-cavern hydrogen storage is rated as high in cross-temporal flexibility because it offers very large-scale and long-duration storage, making it particularly effective for bridging prolonged supply-demand mismatches and supporting resilience during extended disturbances.
+By contrast, hydrogen tanks are rated as medium because they mainly provide short-term to medium-term on-site buffering with more limited storage scale and duration.
+Hydrogen pipelines are also rated as medium, as they can provide a certain degree of temporal balancing through linepack; however, this capability is generally much more limited than that of dedicated geological storage, and pipelines are primarily designed for transport rather than long-duration storage.
+In addition, HTT is mainly characterized by cross-spatial flexibility because its principal function is to transport hydrogen from one location to another.
+Its cross-temporal flexibility is rated as not obvious for two reasons.
+First, the transportation time associated with HTT is essentially a logistics-related delivery lag rather than intentional temporal shifting enabled by storage and deferred use.
+In other words, simply moving hydrogen from one location to another over several hours does not provide the same type of temporal flexibility as storing hydrogen in tanks, pipelines, or salt caverns for later use.
+Second, although HTT may provide limited and scenario-dependent temporal support in practice (e.g., the HTT itself can serve as a form of mobile hydrogen inventory and careful dispatch may allow some degree of deferred delivery under emergency conditions), this effect is indirect, logistics-dependent, and generally weaker than the intrinsic time-shifting capability of dedicated storage technologies.
+It should also be noted that additional resilience benefits can be achieved through the coordinated use of different hydrogen-related technologies.
+For example, combining proton exchange membrane fuel cells with on-site hydrogen tanks can enable fast, clean, and autonomous black start for critical infrastructure such as hospitals.
+Similarly, integrating electrolyzers, salt-cavern storage, pipelines, and hydrogen turbines can support large-scale grid restoration.
+Therefore, careful selection and coordination of hydrogen-related technologies according to specific application scenarios are essential for the design of hydrogen-enabled MESs.
+TABLE III: Resilience Potential of Hydrogen-Related Technologies in MESs (H=High, M=Medium, -=Not obvious)
+Category Technology Cross-temporal flexibility Cross-spatial flexibility Cross-sector flexibility Black-start capability Resilience Potential Explanations
+Hydrogen production Electrolyzer - - H - Converts surplus power to hydrogen; siting is flexible but not mobile; not a generator
+Hydrogen production Steam methane reforming - - H - Centralized, utility-dependent hydrogen production; strong industrial fuel support, but weak emergency agility
+Hydrogen storage Salt-cavern hydrogen storage H - - - Seasonal, very-large storage; immobile; withdrawal needs powered compressors
+Hydrogen storage Hydrogen tank M - - - Hours–days buffer onsite; fixed asset; enables generators but does not generate
+Hydrogen transport Hydrogen pipelines M H - - Pipeline linepack; cross-spatial transport; not a generator
+Hydrogen transport HTT - H - - Mobile hydrogen supply to critical loads; not a generator
+Hydrogen utilization Proton exchange membrane fuel cell - - H H Combined heat and power generation; Fast black-start for critical loads
+Hydrogen utilization Solid oxide fuel cell - - H - Combined heat and power generation; long warm-up; poor for rapid/black-start use
+Hydrogen utilization Hydrogen turbines - - H H Combined heat and power generation; configured units can black-start and support restoration
+II-C Vulnerabilities of Hydrogen-related Technologies
+Although resilience potential can be offered by hydrogen-related technologies in the chain of production, storage, transportation, and utilization, security vulnerabilities also exist in each stage, which can create new disturbance propagation paths and may weaken rather than strengthen system resilience under adverse conditions.
+To be specific, in the production stage, especially in electrolysis and steam methane reforming, the risk of hydrogen leakage and explosion is significant due to high operating pressures and temperatures.
+In the storage stage, compressed hydrogen tanks face threats such as high-pressure rupture, while cryogenic liquid hydrogen presents risks of rapid venting.
 Long-term storage methods like salt caverns must address potential leakage and environmental contamination.
-During the transportation stage, hydrogen’s small molecular size increases the likelihood of leaks in pipelines or tube trailers, and mobile transport systems are vulnerable to traffic accidents or mechanical damage.
+In the transportation stage, hydrogen’s small molecular size increases the likelihood of leaks in pipelines or tube trailers, and mobile transport systems are vulnerable to traffic accidents or mechanical damage.
 Additionally, improper handling during transfer can lead to static ignition or pressure imbalances.
-During the utilization stage, poor ventilation in confined spaces (e.g., tunnels, vehicles, or buildings) can lead to hydrogen accumulation and fire hazards.
-Fuel cell systems may also face risks of overheating or malfunction if not properly managed.
-Across all stages, the growing reliance on automated control and digital monitoring introduces cybersecurity vulnerabilities[59], [60], e.g., hydrogen tank levels and hydrogen demand are vulnerable to false data injection attacks[60].
-These security issues highlight the need for rigorous safety standards, advanced monitoring systems, and hydrogen-compatible materials to ensure safe deployment of hydrogen technologies.
+In the utilization stage, inadequate ventilation may cause hydrogen accumulation and fire hazards.
+Moreover, failures in electrolyzers, hydrogen storage tanks, hydrogen pipelines, or fuel cell systems may compromise local energy supply and even lead to load shedding.
+In addition, the growing digitalization of hydrogen infrastructures introduces cybersecurity threats[59], [60], such as false data injection attacks on tank levels and demand information[60].
+Therefore, hydrogen should be regarded as a technology with both resilience benefits and resilience risks, and its positive contribution to MES resilience depends on proper design, operation, protection, and cybersecurity measures.
 TABLE IV: Optimization Challenges Introduced by Adopting Hydrogen in MES for Resilience Enhancement
-Index Challenges Hydrogen-related Example Formulas / Model Implications Formulations and Optimization Methods
+Index Challenges Hydrogen-related Example Formulas / Model Implications Formulations and Optimization Methods/Techniques
 1 Uncertainties Hydrogen demand varies with mobility or industrial loads $\min_{x\in X}c^{\top}x+\mathbb{E}_{\xi}[Q(x,\xi)]$ Stochastic programming, robust optimization, scenario-based MILP
-2 Mixed integer variables Investment locations of electrolyzers and hydrogen tanks $y_{i}^{\text{electrolyzer}}\in\{0,1\},\quad z_{j}^{\text{tank}}\in\{0,1\}$ MILP, MINLP, branch-and-bound, Lagrangian relaxation
+2 Mixed integer variables Investment locations of electrolyzers and hydrogen tanks $y_{i}^{\text{electrolyzer}}\in\{0,1\},\qquad z_{j}^{\text{tank}}\in\{0,1\}$ MILP, MINLP, branch-and-bound, Lagrangian relaxation
 3 Multi-timescale variables Slow hydrogen storage level update and fast battery dispatch $H_{t}^{\text{stored}}$ (in hours), $B_{\tau}^{\text{stored}}$ (in 5 minutes) MILP, Nested model predictive control, hierarchical deep reinforcement learning
@@
 5 Spatial coupling constraints Hydrogen produced in region A, stored in region B, and used in region C via pipelines $\sum_{j\in\mathcal{N},j\neq i}h_{ji,t}-\sum_{j\in\mathcal{N},j\neq i}h_{ij,t}=h^{\text{net}}_{i,t},~\forall i\in\mathcal{N}$ MILP, Benders decomposition, distributed optimization
-6 Nonlinear & nonconvex constraints Electrolyzer efficiency curve, pressure-flow relationships in pipelines $\eta^{\text{elec}}=f(P_{\text{input}}),\quad Q=k\cdot\sqrt{\Delta P}$ NLP, convex relaxation, piecewise linear approximation
-7 Multi-energy coupling constraints Hydrogen enables electricity-gas-heat-transport sector integration $H_{t}^{\text{prod}}=\eta^{\text{elec}}\cdot P_{t}^{\text{elec}}/\omega^{\text{LHV}}$ Integrated MILP/MINLP, Benders decomposition
-8 Multi-objective function Tradeoff between cost minimization and resilience/load recovery maximization $\min~\text{Cost}$; $\max~\text{Load~Recovery}$ MILP/MINLP, Multi-objective optimization/reinforcement learning
+6 Nonlinear & nonconvex constraints Load-dependent electrolyzer/fuel cell efficiencies, part-load operating regimes, nonlinear hydrogen pipeline hydraulics, compression/liquefaction energy consumption, and nonlinear hydrogen storage state relations $\eta_{t}^{\mathrm{EL}}=f(P_{t}^{\mathrm{EL}})$, $\eta_{t}^{\mathrm{FC}}$ = $g(P_{t}^{\mathrm{FC}})$, $H_{t}^{\mathrm{prod}}$ $\frac{\eta_{t}^{\mathrm{EL}}P_{t}^{\mathrm{EL}}\Delta t}{\omega^{\mathrm{L}}}$, $P_{t}^{\mathrm{FC,out}}$ $\frac{\eta_{t}^{\mathrm{FC}}H_{t}^{\mathrm{FC,in}}\omega^{\mathrm{L}}}{\Delta t}$, $h_{ij,t}$$k_{i,j}(p_{i,t},~p_{j,t})$, $E_{t}^{\mathrm{c}}$$\varphi(q_{t}^{\mathrm{H_{2}}},p_{t}^{\mathrm{in}},p_{t}^{\mathrm{out}})$$p_{t}^{\mathrm{tk}}$$\phi(m_{t}^{\mathrm{H_{2}}},T_{t},V_{t})$ MINLP, NLP, piecewise linear approximation, convex/cone relaxation, sequential linearization, McCormick envelopes, metaheuristics, and deep reinforcement learning
+7 Multi-energy coupling constraints Hydrogen enables electricity-gas-heat-transport sector integration $P_{t}^{\mathrm{FC,e}}=\frac{\eta_{t}^{\text{FC,e}}H_{t}^{\mathrm{FC,in}}\omega^{\mathrm{L}}}{\Delta t}$,  $Q_{t}^{\mathrm{FC,h}}=\frac{\eta_{t}^{\text{FC,h}}H_{t}^{\mathrm{FC,in}}\omega^{\mathrm{L}}}{\Delta t}$ Integrated MILP/MINLP, Benders decomposition
+8 Multi-objective functions Tradeoff between cost minimization and resilience/load recovery maximization $\min~\text{Cost}$; $\max~\text{Load~Recovery}$ MILP/MINLP, Multi-objective optimization/reinforcement learning
 II-D Optimization Challenges Related to Hydrogen-enabled Planning and Operation Enhancement
-Integrating hydrogen into the planning and operation enhancement of MESs introduces a range of complex optimization challenges.
+Integrating hydrogen into the resilience-oriented planning and operation of MESs introduces a range of complex optimization challenges.
 Firstly, uncertainty in hydrogen demand complicates the optimization process, often requiring stochastic or robust optimization frameworks.
@@
 Due to the physical limits on gas flow, thermal dynamics, and control systems, hourly updates of hydrogen storage level are sufficient for modeling storage dynamics.
-In contrast, batteries can be dispatched every 5 minutes to capture the fluctuations of renewable generation and load demand.
-When conducting operation optimization in hours, multi-timescale decision variables are introduced.
+In contrast, batteries could be dispatched every 5 minutes to capture the fluctuations of renewable generation and load demand.
+When conducting operation optimization on an hourly basis, multi-timescale decision variables would arise.
 Fourthly, hydrogen enables cross-temporal operations (e.g., producing and storing energy in low-demand periods for later use), which creates strong inter-temporal coupling and increases the dimensionality of the optimization problem.
 Fifthly, hydrogen transport via pipelines or trailers introduces spatial coupling, requiring network-based flow models with additional constraints.
-Sixthly, several key processes (e.g., electrolyzer efficiency, pressure-flow relationships in pipelines, and fuel cell performance) are inherently nonlinear, leading to nonconvex constraints that complicate solution tractability.
+Sixthly, there are many nonlinear or even nonconvex constraints involved in the hydrogen production and the associated compression, storage, transport, and utilization processes, which complicate solution tractability.
+Taking electrolyzers, fuel cells, hydrogen pipelines, hydrogen compressors/liquefiers, and hydrogen storage tanks into consideration, there are nonlinear or nonconvex constraints incurred by load-dependent electrolyzer/fuel cell efficiencies, part-load operating regimes of electrolyzers/fuel cells, nonlinear hydrogen pipeline hydraulics, energy consumption of hydrogen compressors/liquefiers, and nonlinear hydrogen storage state relations.
 Seventhly, the integration of hydrogen also deepens the coupling between electricity, gas, heat, and mobility sectors, leading to multi-energy coupling constraints.
 Finally, there are multi-objective functions in HMESs, e.g., system cost minimization and operation risk minimization, operation cost minimization and load recovery maximization.
-In summary, the above-mentioned challenges brought by the inclusion of hydrogen call for advanced solution methods, such as mixed-integer nonlinear programming, decomposition techniques, scenario-based planning, robust optimization[61], metaheuristics[60], convex relaxation, model predictive control[62], and deep reinforcement learning[63].
-For the ease of understanding, we summarize the optimization challenges introduced by adopting hydrogen in MES for resilience enhancement in Table IV, where $x$ denotes the first-stage decision variables (e.g., hydrogen storage or electrolyzer sizing); $\xi$ represents the uncertainty parameters (e.g., hydrogen demand, renewable availability); $\Xi$ is the uncertainty set; $c^{\top}x$ is the deterministic first-stage investment cost; $Q(x,\xi)$ is the recourse cost in the second stage given scenario $\xi$.
-$y_{i}^{\text{electrolyzer}}$ is a binary variable indicating whether an electrolyzer is installed at location $i$, $z_{j}^{\text{tank}}$ is a binary variable representing whether a hydrogen storage tank is constructed at node $j$, $y,z=1$ implies the component is built/active, while $y,z=0$ implies it is not.
-$H_{t}^{\text{stored}}$ and $B_{\tau}^{\text{stored}}$ denote the energy level in hydrogen tank and battery at time $t$ and $\tau$, respectively.
-$H_{t}^{\text{prod}}$ and $H_{t}^{\text{used}}$ denote the produced hydrogen and the consumed hydrogen at time $t$, respectively.
-$\mathcal{N}$ is the set of all nodes (e.g., cities, regions, or hydrogen hubs) in the hydrogen network, $\mathcal{T}$ is the set of time periods in the optimization horizon, $h_{ij,t}$ is the amount of hydrogen flow from node $i$ to node $j$ at time $t$, $h_{ji,t}$ is the amount of hydrogen flow from node $j$ to node $i$ at time $t$, $h^{\text{net}}_{i,t}$ is the net hydrogen injection at node $i$ at time $t$, which is positive if hydrogen is injected (e.g., from an electrolyzer) and negative if it is consumed (e.g., by a fuel cell or local demand).
-$\eta^{\text{elec}}$ is the efficiency of the electrolyzer, $P_{\text{input}}$ is the input electrical power, $f(\cdot)$ represents a nonlinear, possibly nonconvex function of power input, $Q$ is the hydrogen flow rate through a pipeline, $\Delta P$ is the pressure difference across the pipeline section, and $k$ is a physical constant related to pipe geometry and gas properties.
-$P_{t}^{\text{elec}}$ is the electrical power supplied to the electrolyzer at time $t$, $\omega^{\text{LHV}}$ denotes the lower heating value of hydrogen.
+It is worth noting that the mathematical models of hydrogen equipment do not necessarily introduce features that are entirely absent from all other MES components, since nonlinearities, mixed-integer decisions, and temporal couplings may also arise in other devices such as CCHP units, batteries, and gas networks.
+However, compared with conventional MES equipment, hydrogen-related devices tend to exhibit these characteristics in a more pronounced and integrated manner, especially when hydrogen production, compression, storage, transport, and utilization are modeled simultaneously.
+In order to address the above challenges, more advanced solution methods should be adopted, such as mixed-integer nonlinear programming, decomposition techniques, scenario-based planning, robust optimization[61], metaheuristics[60], convex relaxation, model predictive control[62], and deep reinforcement learning[63].
+For the ease of understanding, we summarize the optimization challenges introduced by adopting hydrogen in MESs for resilience enhancement in Table IV.
+In line 1, $x$ denotes the first-stage decision variables (e.g., hydrogen storage or electrolyzer sizing); $\xi$ represents the uncertainty parameters (e.g., hydrogen demand, renewable availability); $X$ is the uncertainty set; $c^{\top}x$ is the deterministic first-stage investment cost; $Q(x,\xi)$ is the recourse cost in the second stage given scenario $\xi$.
+In line 2, $y_{i}^{\text{electrolyzer}}$ is a binary variable indicating whether an electrolyzer is installed at location $i$, $z_{j}^{\text{tank}}$ is a binary variable representing whether a hydrogen storage tank is constructed at node $j$, $y,z=1$ implies the component is built/active, while $y,z=0$ implies it is not.
+In line 3, $H_{t}^{\text{stored}}$ and $B_{\tau}^{\text{stored}}$ denote the energy level in hydrogen tanks and batteries at time $t$ and $\tau$, respectively.
+In line 4, $H_{t}^{\text{prod}}$ and $H_{t}^{\text{used}}$ denote the produced hydrogen and the consumed hydrogen at time $t$, respectively.
+In line 5, $\mathcal{N}$ is the set of all nodes (e.g., cities, regions, or hydrogen hubs) in the hydrogen network, $\mathcal{T}$ is the set of time periods in the optimization horizon, $h_{ij,t}$ is the amount of hydrogen flow from node $i$ to node $j$ at time $t$, $h_{ji,t}$ is the amount of hydrogen flow from node $j$ to node $i$ at time $t$, $h^{\text{net}}_{i,t}$ is the net hydrogen injection at node $i$ at time $t$, which is positive if hydrogen is injected (e.g., from an electrolyzer) and negative if it is consumed (e.g., by a fuel cell or local demand).
+In line 6, $\eta_{t}^{\mathrm{EL}}$ and $\eta_{t}^{\mathrm{FC}}$ denote the electrolyzer and fuel cell efficiencies at time $t$, respectively; $f$, $g$, $k_{i,j}$, $\varphi$ and $\phi$ are nonlinear or even nonconvex functions.
+$P_{t}^{\mathrm{EL}}$ and $P_{t}^{\mathrm{FC}}$ denote the electrolyzer input power and fuel cell operating power, respectively; $\Delta t$ is the time-step duration; and $\omega^{\mathrm{L}}$ is the lower heating value of hydrogen.
+$P_{t}^{\mathrm{FC,out}}$ denotes the fuel cell electric output power, and $H_{t}^{\mathrm{FC,in}}$ is the hydrogen input to the fuel cell.
+$p_{i,t}$ and $p_{j,t}$ denote the nodal hydrogen pressures.
+$E_{t}^{\mathrm{c}}$ denotes the compressor energy consumption, $q_{t}^{\mathrm{H_{2}}}$ is the hydrogen flow through the compressor, and $p_{t}^{\mathrm{in}}$ and $p_{t}^{\mathrm{out}}$ denote the inlet and outlet pressures, respectively.
+$p_{t}^{\mathrm{tk}}$ is the hydrogen tank pressure, while $m_{t}^{\mathrm{H_{2}}}$, $T_{t}$, and $V_{t}$ denote the stored hydrogen mass, storage temperature, and tank volume at slot $t$, respectively.
+In line 7, $\eta_{t}^{\mathrm{FC,e}}$ and $\eta_{t}^{\mathrm{FC,h}}$ are the electrical and thermal conversion efficiencies of the fuel cell, respectively.
+$P_{t}^{\mathrm{FC,e}}$ and $Q_{t}^{\mathrm{FC,h}}$ denote the electrical and thermal outputs of the fuel cell, respectively.
 III Hydrogen-enabled Resilience Enhancement Framework for MESs
@@
 As shown in Fig.
-2, the framework of hydrogen-enabled resilience enhancement for MESs is provided, where four steps could be identified, i.e., resilience metrics, event-oriented contingency modeling, planning enhancement, and operation enhancement.
+2, the framework of hydrogen-enabled resilience enhancement for MESs is provided, where four steps could be identified, i.e., resilience metrics, event-oriented contingency modeling and uncertainty processing, planning enhancement, and operation enhancement.
 To be specific, a resilience metric related to the investigated HMESs should be first defined, which would be incorporated into the objectives or constraints of planning or operation problem models.
-Then, according to the type of extreme events (e.g., hurricane, flood, ice disaster, earthquake, cyber attacks), the corresponding event models could be developed, which consist of event-related physical attributes[64], e.g., movement speed and landfall point of typhoons, and the intensity of earthquakes.
+Then, according to the type of extreme events (e.g., hurricane, flood, ice disaster, earthquake, and cyber attacks), the corresponding event models could be developed, which consist of event-related physical attributes[64], e.g., movement speed and landfall point of typhoons, and the intensity of earthquakes.
 Since such attributes depend on some unknown factors, the occurrence of an extreme event is full of uncertainties in the planning stage, e.g., there are many possible landfall scenarios when typhoons are considered[47].
-After obtaining the physical attributes of extreme events, the stochastic impact of extreme events on the contingencies of MESs could be developed.
-Next, taking parameter uncertainty (e.g., renewable generation and multi-type loads), event uncertainty, and event-oriented contingency uncertainty into consideration, scenario generation could be conducted in preparation for planning and operation problems.
+After obtaining the physical attributes of extreme events, the contingencies of MESs incurred by extreme events need to be modeled.
+Typically, stochastic multi-type contingencies are considered in existing works[22].
+In order to cope with the above-mentioned event uncertainties and contingency uncertainties, as well as parameter uncertainty (e.g., renewable generation and multi-type loads), uncertainty processing methods will be adopted, e.g., scenario-based methods, uncertainty-set-based methods, and ambiguity-set-based methods.
+Such methods can provide scenario parameter information, feasible decision space, and probability distribution sets for planning and operation optimization problems.
+More details about uncertainty processing methods and extreme-event scenario generation methods will be introduced in the Table VIII and Table X in Section VI, respectively.
 In the planning stage, some permanent facilities will be deployed, e.g., SHSs, HESSs, hydrogen pipelines, and HRSs.
@@
 In industrial communities, electricity/hydrogen demands can be observed.
-All communities are connected through electricity transportation networks, hydrogen transmission pipelines, and transportation networks.
-With the help of transportation networks, MHERs (e.g., HTTs and HFCVs can be redistributed among multiple communities for improving network resilience, e.g., providing emergent power supply or hydrogen supply to a community when it suffers from an extreme event (e.g., power outages or hydrogen pipeline leakage).
+All communities are connected through electricity distribution networks, hydrogen transmission pipelines, and road transportation networks.
+With the help of transportation networks, MHERs (e.g., HTTs and HFCVs) can be redistributed among multiple communities for improving network resilience, e.g., providing emergent power supply or hydrogen supply to a community when it suffers from an extreme event (e.g., power outages or hydrogen pipeline leakage).
 Figure 3: Typical HMES architecture (building-level)
-Figure 4: Typical HMES architecture (region-level)
+Figure 4: Typical HMES architecture (region-level)[67]
 IV-B Definitions of HMES Resilience
@@
 5, respectively.
-Although the typical curve under MES cannot perfectly capture the realistic system performance trajectories (e.g., local performance fluctuations) as illustrated in [68, 69, 70], it can reflect the system’s ability to prepare for, absorb, and recover from disturbances and has been widely used in existing works[50],[71, 72, 73, 74, 75],[115].
+Although the typical curve under MES cannot perfectly capture the realistic system performance trajectories (e.g., local performance fluctuations) as illustrated in [68, 69, 70], it can reflect the system’s ability to prepare for, absorb, and recover from disturbances and has been widely used in existing works[50],[71, 72, 73, 74, 75, 76].
 In addition, while it is not tailored to hydrogen-enabled MESs, the performance trajectory can be applied to hydrogen-enabled MESs after minor modifications.
@@
 During the extreme event, the emergency response is adopted to minimize the load shedding from time $t_{2}$ to time $t_{3}$ so that the extent of system performance degradation could be alleviated (note that the system performance can be maintained for a short time due to the deployed mobile resources in the preventive response).
-Then, the degraded state could be observed from time $t_{3}$ to time $t_{4}$ at the end of event.
+Then, the degraded state could be observed from time $t_{3}$ to time $t_{4}$ at the end of the event.
 From time $t_{4}$ to time $t_{5}$, the restoration response is achieved to restore critical loads [74] like hospitals and water pumps by adopting temporary measures, such as grid-island operation using remaining storage and mobile energy storage units (e.g., electric vehicles).
 These temporary measures restore partial or prioritized functionality, even if the underlying infrastructure (e.g., transmission lines and distribution equipment) is not yet fully repaired.
-After stabilizing essential services, the system proceeds to recover full infrastructure, e.g., repairing line and equipment.
+After stabilizing essential services, the system proceeds to recover full infrastructure, e.g., repairing lines and equipment.
 Similarly, after integrating hydrogen into MESs, the system performance curve should reflect both the physical and functional characteristics of hydrogen.
 In the preventive stage, additional mobile energy resources (e.g., HTTs for gas delivery and liquid-hydrogen tankers, as well as HFCVs/HFCBs used as mobile generators) can be deployed, so that system performance can be sustained longer in the emergency period.
-During the emergency response, if a grid outage coincides with a hydrogen-pipeline leak, part of critical demand can be supplied by fuel cells.
-Owing to pipeline linepack and pressure regulation, fuel cell stack-inlet pressure/flow conditions can be temporarily maintained, allowing the fuel cell to sustain its set power and thereby reduce load shedding as long as the available upstream pressure/flow remains above the requirements of hydrogen regulator and fuel cell stack.
-If the leak reduces the available supply below those requirements, hydrogen starvation occurs[116], leading to a rapid voltage drop and possible protective shutdown, which increases load shedding.
-Although linepack buffering may delay this adverse impact, fuel-cell power loss can worsen even after the initiating event ends once the buffer is depleted.
+During the emergency response, if a grid outage coincides with hydrogen-pipeline leakage, part of critical demand can be supplied by fuel cells.
+Owing to pipeline linepack and pressure regulation, fuel-cell stack inlet pressure/flow conditions can be temporarily maintained, allowing the fuel cell to sustain its set power and thereby reduce load shedding as long as the available upstream pressure/flow remains above the requirements of hydrogen regulators and fuel cell stacks.
+If the leakage reduces the available supply below those requirements, hydrogen starvation occurs[133], leading to a rapid voltage drop and possible protective shutdown, which increases load shedding.
+Although linepack buffering may delay this adverse impact, fuel-cell power output may continue to decline even after the initiating event ends once the buffer is depleted.
 Since more resources are available for deployment, an HMES can exhibit a higher performance level in the degraded state, and an example of this phenomenon can be seen in Fig.
@@
 Utilizing mobile resources or conducting pipeline reconfiguration, the recovery trajectory of HMES can also be raised.
-Compared with MESs, hydrogen-related equipment failure in HMESs may require longer restoration time as shown in Fig.
-5, which rely on random transportation network state and hydrogen flow dynamics as shown in Fig.
-6 of [120] in the revised manuscript.
+Compared with MESs, hydrogen-related equipment failures in HMESs may involve longer restoration time, as illustrated in Fig.
+5.
+This extended recovery is influenced by stochastic transportation network states and hydrogen flow dynamics as supported by Fig.
+6 of [135].
 Based on the above description, HMES resilience can be defined as the capacity of an HMES to withstand the disturbance from an extreme event (i.e., event progress during $t_{2}$ and $t_{3}$), and to rebuild and renew it afterwards[9].
 Similarly, HMES resilience can also be described as the capacity to limit the system performance degradation and the duration of the degraded state in order to maintain critical services following an extreme event[5].
-In addition to energy resilience[118], carbon resilience is also proposed in existing works, which describes an energy system’s ability to maintain or recover to its current and predicted emission level during and after shocks such as renewable intermittency, public health emergencies, or climate-related disruptions[53].
 Although many definitions related to HMES resilience are proposed, they are originally borrowed from power grids and MESs, which can not truly reflect the physical and functional features of hydrogen[18][28][71, 72, 73, 74, 75].
-For example, when an extreme event (e.g., ice disaster) causes simultaneous power outages and hydrogen pipeline leakage, even if no load loss is incurred due to backup sources, hydrogen fuel cells can not operate normally due to the loss of pressurized hydrogen supply.
+For example, when an extreme event (e.g., ice disaster) causes simultaneous power outages and hydrogen pipeline leakage, even if no load loss is incurred due to backup sources, hydrogen fuel cells may not operate normally due to the loss of pressurized hydrogen supply.
 At this time, no resilience enhancement measure is required since there is no system performance degradation (i.e., load loss).
 In fact, hydrogen-related systems need immediate repair so that further pressure reduction can be avoided.
-Otherwise, once backup sources are unavailable, the whole HMES may collapse and result in larger negative social-economical impacts.
-Therefore, the definition of HMES resilience should incorporate any negative impacts (e.g., hydrogen availability and hydrogen-demand loss) of extreme events on hydrogen energy resources into consideration.
+Otherwise, once backup sources are unavailable, the whole HMES may collapse and result in larger negative social and economic impacts.
+Therefore, the definition of HMES resilience should take any negative impacts (e.g., hydrogen availability and hydrogen-demand loss) of extreme events on hydrogen energy resources into consideration.
 IV-C Metrics of HMES Resilience
@@
 where $t_{2}$ and $t_{5}$ denote the event happening time and full recovery time as shown in Fig.
-4, $F^{\min}$ denotes the minimum system performance requirement, and $F(t)>F^{\min}]=1$ if $F(t)>F^{\min}$.
+3, $F^{\min}$ denotes the minimum system performance requirement, and $F(t)>F^{\min}]=1$ if $F(t)>F^{\min}$.
 Otherwise, $F(t)>F^{\min}]=0$.
@@
 Therefore, to develop a comprehensive resilience metric, it is necessary to combine load-based and customer-based metrics, as mentioned in [93].
-In addition, existing metrics can not reflect the degraded state (e.g., HESS failure or hydrogen pipeline leaking) of HERs during an extreme event.
+In addition, existing metrics can not reflect the degraded state (e.g., HESS failure or hydrogen pipeline leakage) of HERs during an extreme event.
 Therefore, more comprehensive metrics are expected.
@@
 $\Omega=\{\boldsymbol{u}|\sum\nolimits_{e}u_{\text{EL},e}+\sum\nolimits_{f}u_{\text{FC},f}+u_{\text{GT}}\geq N_{g}\},$ (15)
-where $\boldsymbol{u}=u_{\text{EL},e},~u_{\text{FC},e},~u_{\text{GT}}$.
+where $\boldsymbol{u}=(u_{\text{EL},e},~u_{\text{FC},e},~u_{\text{GT}})$.
 Moreover, $u_{\text{EL},e}$$u_{\text{FC},e}$$u_{\text{GT}}$ are binary variables related to electrolyzers, fuel cells, and a gas turbine, respectively.
@@
 In addition, some cascading contingencies incurred by cyber-physical attacks are not considered in existing works[103].
-Above all, existing works rarely consider hydrogen-specific vulnerabilities.
+Overall, existing works rarely consider hydrogen-specific vulnerabilities.
 In fact, there are many physical vulnerabilities and cyber vulnerabilities in HMESs.
-In physical aspects, hydrogen tends to cause material embrittlement, particularly in metals such as steel, can lead to long-term degradation of pipelines, storage tanks, and connectors.
+In physical aspects, hydrogen tends to cause material embrittlement, particularly in metals such as steel, which can lead to long-term degradation of pipelines, storage tanks, and connectors.
 These structural vulnerabilities increase the likelihood of mechanical failure under normal operations or stress conditions such as temperature fluctuations and pressure surges during extreme events.
@@
 However, most contingency models treat hydrogen assets as idealized components with fixed performance characteristics in existing works, ignoring the risks posed by leakage, material fatigue, and cyber intrusion.
-Therefore, future research should develop risk-aware, vulnerability-inclusive contingency modeling that capture both the physical fragilities and cyber risks related to hydrogen.
+Therefore, future research should develop risk-aware, vulnerability-inclusive contingency modeling that captures both the physical fragilities and cyber risks related to hydrogen.
 VI Hydrogen-enabled Planning Enhancement for MESs
@@
 Simulation results showed that the proposed planning method can reduce the total cost compared to other four baselines by 6.36%-50.54%, which focus on only resilience (load shedding), only investment economics, only HRS planning, and only gas-fired distributed generators.
-Figure 7: The framework of two-stage resilience planning method
+Figure 7: The framework of two-stage resilience planning method[97]
 TABLE VI: Summary of Planning Enhancement for HMES (Single-type Facility)
-Ref. Object Events and contingencies Hydrogen facility investment decisions Demand types Parameter/event uncertainty Optimization objectives Problem types, solving methods
-[103] Power systems based on HESSs Deterministic impacts caused by heatwave events HESS (capacity) Electricity $\times$ Minimizing the total cost Single-level deterministic MILP, CPLEX
-[28] Hydrogen-integrated airport energy system Insufficient power supply capacity caused by extreme events HESS (capacity) Electricity, hydrogen $\times$ Minimizing the overall economic cost Single-level deterministic MILP, CPLEX
-[96] Integrated electricity-gas energy system with HRSs Stochastic line outages incurred by typhoons HRS (capacity, location, quantity) Electricity, gas $\times$ Minimizing the sum of investment cost and expected operation cost, minimizing operation costs under scenarios Two-stage MISOCP, Gurobi
-[104] Hydrogen microgrids Deterministic line outages by natural disasters HESS (capacity, location, quantity) Electricity $\times$ Minimize the total cost; maximize the operation cost by selecting the worst extreme scenario; minimize load shedding and switch action costs Tri-level robust MILP, mixed-integer L-shaped method
-[46] Electricity-hydrogen island energy system Stochastic multi-device failures and power interruptions during extreme events HESS (capacity) Electricity, hydrogen Wind power and electric load Maximizing the total profit, minimum value in the third level, maximizing profit minus load shedding cost Tri-level robust MILP, nested C&CG
-[106] European electricity systems Low renewable generation during Dunkelflaute events HESS (location, capacity) Electricity Renewable generation Minimize total annualized costs; maximizes the operation cost by selecting the worst scenario; minimize the operation cost after the uncertainty realization Tri-level robust MILP; C&CG
-[47] Electricity-hydrogen integrated energy system Stochastic line outages incurred by typhoons H2P (capacity, location, quantity) Electricity, hydrogen Event uncertainty Minimizing investment cost and expected worst-case system cost Tri-level stochastic robust optimization problem, nested C&CG with progressive hedging
-[107] Gas-electric-hydrogen microgrids NkN-k contingencies caused by extreme events HESS (location, quantity) Electricity, heat $\times$ Minimizing the total cost, maximizing load restoration Bi-level MILP, CPLEX
+Ref. Parameter/event uncertainty Contingency uncertainty Uncertainty processing methods Hydrogen facility investment decisions Optimization objective components Problem types and solving methods
+[104] $\times$ Deterministic impacts $\times$ HESS (capacity) Total cost Single-level deterministic MILP, CPLEX
+[28] $\times$ Deterministic impacts $\times$ HESS (capacity) Overall economic cost Single-level deterministic MILP, CPLEX
+[97] $\times$ Line fault uncertainty Line fault scenarios HRS (capacity, location, quantity) Investment cost and expected operation cost Two-stage MISOCP, Gurobi
+[105] Natural disaster event uncertainty Component faults and fault durations Uncertainty set, natural disaster scenarios HESS (capacity, location, quantity) Total cost, operation cost, load shedding, and switching cost Tri-level robust MILP, mixed-integer L-shaped method
+[46] $\times$ Fault seasons and component fault combinations Season scenarios, fault uncertainty sets HESS (capacity) Total profit, load shedding cost Tri-level robust MILP, nested C&CG
+[107] $\times$ Low renewable generation Cardinality-constrained uncertainty sets HESS (location, capacity) Total annualized costs; operation cost Tri-level robust MILP; C&CG
+[47] Typhoon event uncertainty Stochastic line outages incurred by typhoons Typhoon scenarios, line outage uncertainty sets H2P (capacity, location, quantity) Investment cost and expected worst-case system cost Tri-level stochastic robust optimization, nested C&CG with hedging
+[108] $\times$ NkN-k contingencies Contingency scenarios HESS (location, quantity) Total cost, load restoration quantity Bi-level MILP, CPLEX
 Figure 8: The framework of the tri-level robust planning method
@@
 Simulation results indicated the importance of considering widespread regional weather events when designing a robust Europe electricity system.
+Figure 9: Illustration of line outages in an electricity-hydrogen integrated energy system under typhoons[47]
 Although stochastic effects of an event have been considered in the above-mentioned works, they neglect the event uncertainty.
@@
 Simulation results showed that the proposed approach can enhance the system’s resilience by the adoption of the microgrid and reduce expected electric load loss by 30%-34%.
-Figure 9: Illustration of line outages in an electricity-hydrogen integrated energy system under typhoons
 TABLE VII: Summary of Planning Enhancement for HMES (Multi-type facilities)
-Ref. Object Events and contingencies Hydrogen facility investment decisions Demand types Parameter/event uncertainty Optimization objectives Problem types, solving methods
-[108] Integrated gas and electricity systems with HESSs Two kinds of deterministic extreme weather events HESS, hydrogen production facilities (capacity, location) Electricity, heating $\times$ Minimizing the total cost Deterministic single-level MINLP, Xpress solver
-[99] Hydrogen-based microgrids NkN-k stochastic device failure caused by disastrous events HESS, hydrogen-driven CCHP, SHS (capacity, location, quantity) Electricity, heating, cooling Renewable generation and demand Minimizing the total cost, minimizing the operation cost Two-stage stochastic MINLP, dual cutting-plane decomposition algorithm
-[109] Hydrogen-based seaport MESs NkN-k line outage contingencies Fuel cell electric truck, HRS, renewable generator, and remote control control (capacity, location) Electricity, heating, cooling Renewable generation, demand, and price Minimizing investment cost and expected operation cost Two-stage stochastic MINLP, Gurobi
-[110] Hydrogen-enriched power distribution networks NkN-k contingencies caused by extreme events HESS, MHER (location, capacity) Electricity, hydrogen Renewable generation, electrical and hydrogen load, and network contingencies Minimize the total cost; minimize the expected operation cost Risk-averse two-stage stochastic MILP, Gurobi
-[22] Hydrogen-based microgrids and hydrogen fuel cell bus NkN-k line contingencies incurred by extreme weather HRS and hydrogen fuel cell electric bus (capacity, location, quantity) Electricity, hydrogen $\times$ Maximizing profit, minimum value of maximized load restoration, maximizing load restoration Tri-level MILP, nested C&CG
-[111] Electricity-hydrogen distribution networks Stochastic power line and hydrogen pipeline contingencies caused by extreme weather Hydrogen pipeline hardening, HESS (capacity) Electricity, hydrogen $\times$ Minimizing the maximum load shedding under the worst-case scenario, minimizing load shedding Distributionally robust MISOCP, C&CG algorithm
-[112] Hydrogen-electrical microgrids for sea islands NkN-k contingencies caused by extreme wind events HESS (location, capacity), hydrogen vessel purchasing and routing Electricity Electricity demand, renewable generation, line outages, hydrogen vessel availability Minimize total annualized costs; minimize operational and maintenance costs under the worst-case probability distribution Distributionally robust MILP, C&CG with strong cutting planes
-[113] Renewable-dominated power systems Stochastic line and tower damage under extreme weather Fuel cell, battery, electrolyzer, and hydrogen tank (capacity, rated power, location) Electricity, hydrogen Renewable generation and demand Minimizing configuration cost and operation cost Bi-level problem, PSO and MILP
-TABLE VIII: Modeling Frameworks of Planning Enhancement Problem for HMESs
+Ref. Parameter/event uncertainty Contingency uncertainty Uncertainty processing methods Hydrogen facility investment decisions Optimization objective components Problem types and solving methods
+[109] $\times$ Deterministic impacts $\times$ HESS, hydrogen production facilities (capacity, location) Total cost Single-level deterministic MINLP, Xpress solver
+[100] Renewable generation and demand NkN-k stochastic device failure Normal operation scenarios, emergency operational scenarios HESS, hydrogen-driven CCHP, SHS (capacity, location, quantity) Total cost, operation cost Two-stage stochastic MINLP, dual cutting-plane algorithm
+[110] Renewable generation, demand, and price NkN-k line outage contingencies Normal operation scenarios, emergency operational scenarios Fuel cell electric truck, HRS, renewable generator, and remote switch (capacity, location) Investment cost, expected operation cost Two-stage stochastic MINLP, Gurobi
+[111] Renewable generation, electrical and hydrogen load NkN-k contingencies Normal operation scenarios, network disruption scenarios HESS, MHER (location, capacity) Total cost, expected operation cost Risk-averse two-stage stochastic MILP, Gurobi
+[22] $\times$ NkN-k line contingencies Cardinality-constrained uncertainty sets HRS and hydrogen fuel cell electric bus (capacity, location, quantity) Profit, load restoration Tri-level MILP, nested C&CG
+[112] Wind speed and rainfall intensity Power line and hydrogen pipeline uncertainties Endogenous second-order moment ambiguity set, scenarios Hydrogen pipeline hardening, HESS (capacity) Load shedding Distributionally robust MISOCP, C&CG algorithm
+[113] Demand, renewable generation, line outages, vessel availability NkN-k contingencies Decision dependent-based ambiguity set, scenarios HESS (location, capacity), hydrogen vessel purchasing and routing Total costs; operational and maintenance costs Distributionally robust MILP, C&CG with strong cutting planes
+[114] $\times$ Deterministic impacts $\times$ Fuel cell, battery, electrolyzer, and hydrogen tank (capacity, rated power, location) Configuration cost, operation cost Bi-level problem, PSO and MILP
+TABLE VIII: Classification of Uncertainty Processing Methods in HMES Planning Studies
+Category Key idea Typical representations Representative characteristics Formulation framework
+Scenario-based methods Uncertainty is represented by a finite number of discrete scenarios, each corresponding to a possible realization of uncertain parameters, extreme events, or contingencies.
+Renewable generation scenarios[100]; load scenarios[110]; natural disaster scenarios[105]; line outage scenarios[47] Suitable for explicitly modeling the evolution of uncertain events; may suffer from scenario explosion and high computational burden Stochastic programming
+Uncertainty-set-based methods Uncertainty is characterized by deterministic sets, and decisions are optimized to remain feasible or robust against all realizations within the prescribed sets.
+Interval sets; box sets; cardinality-constrained uncertainty sets[22][107] Does not require exact probability distributions; usually provides strong robustness guarantees, but may be conservative Robust optimization
+Ambiguity-set-based methods Uncertainty is modeled through a family of probability distributions rather than a single known distribution, and the decision is optimized against the worst-case distribution within the ambiguity set.
+Moment-based ambiguity sets[112]; decision-dependent ambiguity sets[113] Suitable when the exact probability distribution is unknown but partial statistical information is available; balances robustness and probabilistic information Distributionally robust optimization
+TABLE IX: Modeling Frameworks and Solving Methods of Planning Enhancement Problem for HMESs
 Modeling framework Model explanation Mathematical formulation Solution method
-Single-level deterministic optimization Classical optimization with known parameters; no uncertainty or hierarchical decision structure $\min_{x\in X}c^{\top}x+f(x)$ Linear/Nonlinear/Mixed-integer programming solvers (e.g., Gurobi, CPLEX, IPOPT, Xpress)[103][108][28]
-Two-stage stochastic programming First-stage “here-and-now” decisions made before uncertainty is realized; second-stage “wait-and-see” recourse decisions after uncertainty is revealed $\min_{x\in X}c^{\top}x+\mathbb{E}_{\xi}[Q(x,\xi)]$, where $Q(x,\xi)=\min_{y}\{q^{\top}y:Wy+Tx\geq h\}$ Column-and-Constraint Generation (C&CG)[111], dual cutting-plane decomposition[99], commercial solvers (e.g., Gurobi[109][110][96])
+Single-level deterministic optimization Classical optimization with known parameters; no uncertainty or hierarchical decision structure $\min_{x\in X}c^{\top}x+f(x)$ Linear/Nonlinear/Mixed-integer programming solvers (e.g., Gurobi, CPLEX, IPOPT, Xpress)[104]
+Two-stage stochastic programming First-stage decisions made before uncertainty is realized; second-stage recourse decisions made after uncertainty is revealed $\min_{x\in X}c^{\top}x+\mathbb{E}_{\xi}[Q(x,\xi)]$, where $Q(x,\xi)=\min_{y}\{q^{\top}y:Wy+Tx\geq h\}$ Column-and-Constraint Generation (C&CG)[112], dual cutting-plane decomposition[100], commercial solvers (e.g., Gurobi[110][111])
 Tri-level robust optimization Hierarchical decision structure with robustness against worst-case uncertainty $\min_{x\in X}c^{\top}x+\max_{\xi\in\mathcal{U}}\min_{y\in Y(x,\xi)}f(x,y,\xi)$ L-shaped method[105], nested C&CG[46][22], C&CG[107]
-Tri-level stochastic robust optimization Combines stochastic and robust elements: outer stage handles here-and-now, middle stage captures uncertainty distribution/worst-case, inner stage executes corrective recourse $\min_{x}c^{\top}x+\mathbb{E}_{\xi}[\max_{u\in\mathcal{U}(\xi)}\min_{y}f(x,y,u)]$ Nested C&CG with progressive hedging[47]
-Distributionally robust optimization Decisions are robust against the worst-case probability distribution within an ambiguity set, rather than a single known distribution $\min_{x\in X}c^{\top}x+\sup_{P\in\mathcal{P}}\mathbb{E}_{\xi\sim P}[f(x,\xi)]$, where $f(x,\xi)=\min_{y\in Y(x,\xi)}\{q^{\top}y:Wy+Tx\geq h\}$ C&CG with strong cutting planes[112]
+Tri-level stochastic robust optimization Outer stage handles here-and-now, middle stage captures uncertainty distribution/worst-case, inner stage executes corrective recourse $\min_{x}c^{\top}x+\mathbb{E}_{\xi}[\max_{u\in\mathcal{U}(\xi)}\min_{y}f(x,y,u)]$ Nested C&CG with progressive hedging[47]
+Distributionally robust optimization Decisions are robust against the worst-case probability distribution within an ambiguity set, rather than a single known distribution $\min_{x\in X}c^{\top}x+\sup_{P\in\mathcal{P}}\mathbb{E}_{\xi\sim P}[f(x,\xi)]$ s.t.
+$f(x,\xi)=\min_{y\in Y(x,\xi)}\{q^{\top}y:Wy+Tx\geq h\}$ C&CG with strong cutting planes[113]
 Bi-level optimization Leader–follower structure where the upper level (leader) decision depends on the optimal response of the lower level (follower) $\min_{x\in X}c^{\top}x+F(x,y)$ s.t.
@@
 In [114], Wang et al.
-developed a resilient planning method for hybrid hydrogen and battery energy storage systems under extreme events based on particle swarm optimization and MILP, where particle swarm optimization is adopted to decide the system configuration (including installed location of fuel cell and battery; the rated power of fuel cell, battery, electrolyzer, and hydrogen tank; the capacity of battery) and the MILP related to the system operation cost minimization is solved by commercial solvers.
+developed a resilient planning method for hybrid hydrogen and battery energy storage systems under extreme events based on particle swarm optimization and MILP, where particle swarm optimization is adopted to decide the system configuration (including installed location of fuel cells and batteries; the rated power of fuel cells, batteries, electrolyzers, and hydrogen tanks; the capacity of batteries) and the MILP related to the system operation cost minimization is solved by commercial solvers.
 In this work, parameter uncertainty and event uncertainty are not considered, while contingency modeling related to renewable energy reduction, load increase, and line capacity reduction is considered.
 Simulation results indicated that the hybrid energy system can improve the resilience index (i.e., load loss ratio) by 23.8% and 0.7%, compared with single battery energy storage or hydrogen energy storage, respectively.
-Figure 10: The framework of the two-stage planning method
+Figure 10: The framework of the two-stage planning method[100]
 Summary: To show the differences of existing works more clearly, their main features are summarized in Table VI and Table VII.
-It can be observed that not all works considered three kinds of uncertainties simultaneously, e.g., extreme event uncertainty, event-incurred contingency uncertainty, and system parameter uncertainty.
-Moreover, when more complex HMESs are considered (e.g., electricity-hydrogen-gas systems), the number of decisions related to different components becomes larger.
-Consequently, the formulated problems have more complex forms, e.g., MISOCP and MINLP.
-In addition, there are six kinds of modeling frameworks for planning enhancement problem of HMESs, i.e., single-level deterministic programming, bi-level programming, two-stage stochastic programming, three-level robust programming, three-level stochastic robust programming, and three-level distributionally robust optimization.
-Their modeling explanations, mathematical formulation, and solution method are summarized in Table 8.
+It can be observed that there are typically three kinds of uncertainty processing methods in existing works.
+To better understand the differences among these methods, we provide a clear classification of them, as shown in Table VIII.
+Among them, scenario-based methods are suitable for explicitly modeling the evolution of uncertain events, and related planning problems typically follow the framework of stochastic programming.
+Instead of representing uncertainty by a finite number of scenarios, uncertainty-set-based methods adopt deterministic sets to describe uncertainties, e.g., interval sets and cardinality-constrained uncertainty sets.
+These methods do not require exact probability distributions, and related planning problems typically follow the robust optimization or stochastic robust optimization framework.
+When the exact probability distribution is unknown but partial statistical information is available, ambiguity-set-based methods can be adopted, which model uncertainty through a family of probability distributions.
+In this case, related planning problems typically follow the framework of distributionally robust optimization.
+Moreover, when more complex HMESs are considered (e.g., electricity-hydrogen-gas systems), the number of decisions related to different components becomes larger and more nonlinear or nonconvex constraints incurred by hydrogen equipment (e.g., electrolyzers, fuel cells, hydrogen pipelines, hydrogen compressors/liquefiers, and hydrogen storage tanks) are introduced.
+Consequently, the formulated problems have more complex forms (e.g., a MINLP problem is incurred due to the nonlinear hydrogen pipeline hydraulics (i.e., equation (28) in [109])) and advanced solution methods are expected.
+Furthermore, there are six kinds of modeling frameworks for planning enhancement problem of HMESs and we summarize their differences in Table IX.
 In row 1, $x$ denotes all decision variables, $c^{\top}x$ denotes the investment cost, $f(x)$ denotes other cost except investment cost; In row 2, $y$ denotes decision variables of lower-level optimization problem with objective function $f(x,y)$, $F(x,y)$ denotes the operation cost in the upper-level optimization problem; In row 3, $Q(x,\xi)$ denotes the operation cost under the scenario $\xi$; In row 4, $f(x,y,\xi)$ denotes the operation cost under the scenario $\xi$; In row 5, $f(x,y,u)$ denotes the operation cost under a decision $u$, which belongs to $\mathcal{U}(\xi)$; In row 6, $p$ denotes a probability distribution, $f(x,\xi)$ denotes the optimal operation cost under the scenario $\xi$ and decision $x$.
 Finally, the above works mainly focus on simultaneous line outages and/or device faults incurred by natural extreme events, while cascading faults and cyber attacks are seldom considered[115].
+In addition, a few of them considered scenario generation for extreme events based on generative adversarial networks[47] and natural disaster models[105].
+Although such methods have some advantages, they face some limitations in aspects of physical implausibility and bias correction.
+To show the advantages and disadvantages of other scenario generation methods, we provide a classification of extreme-event scenario generation methods in Table X, where representative methods, key idea, advantages, and limitations could be observed.
+TABLE X: Classification of extreme-event scenario generation methods considering the uncertainty of extreme events themselves
+Category Representative methods Key idea Advantages Limitations
+Statistical sampling-based scenario generation Historical event selection[116]; bootstrap resampling[117]; block bootstrapping[118]; parametric distribution fitting with random sampling[47]; representative-period extraction[119] Generate extreme-event scenarios directly from historical observations or their statistical properties, such as empirical frequencies, fitted marginal distributions, or resampling rules Simple and transparent; easy to implement; low computational burden; easy to integrate into stochastic planning models; preserves some historical temporal characteristics Difficult to generate unprecedented or very rare events; limited ability to represent compound extremes and high-dimensional dependence
+Dependence-aware probabilistic modeling Copula-based models[120][121]; multivariate parametric distributions[122]; Bayesian networks[123]; spatio-temporal probability modeling[124] Construct explicit joint probabilistic relationships among multiple variables, so that generated scenarios preserve inter-variable, spatial, and temporal dependence structures Better captures spatial correlation and temporal clustering; more suitable than simple resampling for multi-variable extreme-event modeling Model estimation becomes difficult in high-dimensional settings; tail dependence is hard to identify under limited extreme samples
+Climate- and weather-model-driven scenario generation Regional climate models[105]; global climate model outputs[117]; numerical weather prediction models[125]; bias-corrected climate projections[126]; weather generators[127] Generate extreme-event scenarios from physically based simulation or projection of meteorological and climate processes, rather than only from historical observations Physically grounded; suitable for future climate conditions and long-return-period events; capable of representing persistent and evolving meteorological extremes Requires downscaling, bias correction, and large computational effort; outputs may contain structural uncertainty from the underlying climate/weather models
+Artificial intelligence-enabled synthetic scenario generation Generative adversarial networks[47][128][125]; variational autoencoders[120]; diffusion models[129][130]; hybrid deep generative models[131] Learn complex nonlinear spatio-temporal patterns from data and generate synthetic extreme-event trajectories beyond direct historical resampling Flexible in handling high-dimensional and nonlinear patterns; potentially better at capturing complex spatio-temporal structures than classical approaches Usually data-hungry; may generate physically implausible scenarios without domain constraints; interpretability is weak
 VII Hydrogen-enabled Operation Enhancement for MESs
@@
 proposed a robust hydrogen-penetrated formation strategy against emergent power outages caused by extreme events with the help of HRSs and hydrogen transit while taking uncertainties in load and hydrogen demands, renewable generators, and temporal distances between HRSs into consideration.
+Figure 11: The key idea of the proposed scheduling method for the hydrogen-based multi-energy supply microgrid[18]
 […]
-Figure 11: The key idea of the proposed scheduling method for the hydrogen-based multi-energy supply microgrid
+TABLE XI: Summary of Operation Enhancement for HMES (Single-stage)
+Ref. Parameter/event uncertainty Contingency uncertainty Uncertainty processing methods Operation stages Problem types and solving methods Measure types
+[86] $\times$ Deterministic grid outage $\times$ Preventive MILP; CPLEX A
+[85] $\times$ Deterministic grid outage $\times$ Preventive MINLP; Generalized benders decomposition A
+[95] $\times$ Deterministic line and generator outages $\times$ Preventive MINLP A
+[79] Wind speed, solar radiation, active and reactive loads Deterministic grid outage Scenarios Preventive Stochastic MILP, fuzzy method A
+[29] Demands and renewable generation Deterministic grid outage Scenarios Preventive Stochastic MILP; CPLEX A
+[30] Demands, renewable generators, HRS distances Deterministic grid outage Scenarios Preventive Stochastic MILP; Gurobi B
+[51] Renewable energy and electricity prices Deterministic grid outage Scenarios Emergency Stochastic MILP; CPLEX D
+[137] $\times$ Deterministic grid outage $\times$ Emergency Fuzzy multi-objective MISOCP; Gurobi D
+[134] $\times$ Deterministic line outages and generator trips $\times$ Emergency MILP, CPLEX E
+[31] $\times$ Deterministic grid outage and gas shortage $\times$ Emergency MILP; Alternating direction method of multipliers E
+[18] $\times$ Deterministic component failures $\times$ Emergency MILP, Frank-Wolfe algorithm F
+[76] Loads, renewables, electric vehicle charging behaviors Deterministic component outages Scenarios Emergency Stochastic MILP, GAMS solver H
+[136] PV and wind turbine generation Extreme PV and wind turbine generation Scenarios Emergency Stochastic MILP, Gurobi/CPLEX E, G
+[59] $\times$ Cyber-attacks on demand $\times$ Emergency MILP; CPLEX G
+[80] $\times$ Deterministic line and pipeline outages $\times$ Restoration MISOCP; Relaxation inducement K
+[74] Load demand Line outages Scenarios, ambiguity sets Restoration Chance-constrained adaptive distributionally robust MILP; CPLEX K
+[57] $\times$ Deterministic large-scale blackouts $\times$ Restoration MILP L
+[81] Wind forecast Deterministic line outages Scenarios Restoration Stochastic MILP; Gurobi I
+[78] Renewable outputs, electric and thermal loads Deterministic line outages Scenarios Restoration Stochastic MILP; Gurobi J
+[135] $\times$ Power line and pipeline outages $\times$ Restoration MILP, Gurobi M
 Figure 12: The illustration of an energy hub under threats of false data injection attacking
+Figure 13: The framework of the two-level restoration strategy for a distribution network with HRSs[78]
 After the extreme events, some measures related to system topologies or resources could be adopted to restore load.
@@
 Different from above works, Wang et al.[135] proposed a restoration strategy for an integrated power and hydrogen distribution system with the consideration of power line and hydrogen pipeline outages caused by extreme events.
-When power outages happen, transmission line and hydrogen pipeline reconfiguration was first adopted to avoid islanding.
-Then, repair crews were dispatched to repair the fault lines and hydrogen pipelines.
-In addition, mobile battery-carried vehicles were dispatched to cooperate with repair crews to restore critical loads.
-Figure 13: The framework of the two-level restoration strategy for a distribution network with HRSs
+When power outages occur, reconfigurations of transmission line and hydrogen pipeline are first adopted to avoid islanding.
+Then, repair crews are dispatched to repair the fault lines and hydrogen pipelines.
+In addition, mobile battery-carried vehicles are dispatched to cooperate with repair crews to restore critical loads.
+Figure 14: The framework of the rolling dispatch method with an alert mechanism[83]
 VII-B Multi-stage operation strategy
@@
 Once the event ending is detected, the alert scheduling stage will transfer to the normal scheduling stage.
-Since the function of alert scheduling is similar to that of preventive response (i.e., making preparation for emergency response), their roles are regarded as equivalent in Table VI.
+Since the function of alert scheduling is similar to that of preventive response (i.e., making preparation for emergency response), their roles are regarded as equivalent in Table XI.
 Similar rolling horizon optimization-based methods considering economic operation, proactive preparation, and emergency operation can be observed in [90].
-Figure 14: The framework of the rolling dispatch method with an alert mechanism
-TABLE IX: Summary of Operation Enhancement for HMES (Single-stage)
-Ref. Events Uncertainties Demand types Operation stages Problem types and solving methods Measure types
-[85] Main grid power outages caused by extreme events $\times$ Electricity Preventive MILP; CPLEX Type-A
-[84] Main outages caused by extreme events $\times$ Electricity Preventive MINLP; Generalized Benders Decomposition Type-A
-[94] Line outages and generator disruption $\times$ Electricity, hydrogen Preventive MINLP Type-A
-[78] Main grid power outages caused by extreme events Wind speed, solar radiation, active and reactive loads Electricity, hydrogen Preventive Stochastic MILP, fuzzy method Type-A
-[29] Main grid power outages caused by natural disaster Active and reactive demands and renewable generation Electricity, hydrogen Preventive Stochastic MILP; CPLEX Type-A
-[30] Main grid power outages caused by extreme events Electricity and hydrogen demands, renewable generators, temporal distances between HRSs Electricity, hydrogen Preventive Stochastic MILP; Gurobi Type-B
-[51] Main grid power outages Renewable energy and electricity prices Electricity, hydrogen Emergency Stochastic MILP, CPLEX Type-D
-[121] Main grid power outages caused by extreme events $\times$ Electricity, hydrogen Emergency Fuzzy multi-objective MISOCP; Gurobi Type-D
-[117] Line outages and generator trips caused by ice disasters Line outage probability Electricity, hydrogen Emergency MILP, CPLEX Type-E
-[31] Main grid power outage and insufficient natural gas supply caused by extreme events $\times$ Electricity, gas Emergency MILP; Alternating direction method of multipliers Type-E
-[18] Component failures caused by hurricanes $\times$ Electricity, gas, heat Emergency MILP, Frank-Wolfe algorithm Type-F
-[115] Component outage caused by extreme events Loads, renewables, electric vehicle charging behaviors Electricity, heat, cooling Emergency Stochastic MILP, GAMS solver Type-H
-[119] Extreme PV and wind turbine generation PV and wind turbine generation Electricity, heat, hydrogen Emergency Stochastic MILP, Gurobi/CPLEX Type-E, Type-G
-[59] False data injection cyber-attacks on electricity and hydrogen demand $\times$ Electricity, heat, hydrogen Emergency MILP; CPLEX Type-G
-[79] Distribution line and pipeline outages caused by extreme events $\times$ Electricity, heat, hydrogen Restoration MISOCP; Relaxation inducement Type-K
-[74] Line outages caused by natural disasters Load demand Electricity Restoration Chance-constrained adaptive distributionally robust MILP; CPLEX Type-K
-[57] Large-scale blackouts caused by extreme events $\times$ Electricity Restoration MILP Type-L
-[80] Line outages caused by extreme weather events Wind forecast Electricity Restoration Stochastic MILP; Gurobi Type-I
-[76] Line outages caused by extreme events Renewable outputs, electric and thermal loads Electricity, heat, hydrogen Restoration Stochastic MILP; Gurobi Type-J
-[120] Power line and hydrogen pipeline outages caused by extreme events $\times$ Electricity, hydrogen Restoration MILP, Gurobi Type-M
-TABLE X: Summary of Operation Enhancement for HMES (Multi-stage)
-Ref. Events Uncertainties Demand types Operation stages Problem types and solving methods Hydrogen-related measure types
-[86] Critical component contingencies caused by extreme events Network contingencies Electricity, heat Preventive, emergency Stochastic MISOCP; Receding horizon method Type-A, Type-D
-[82] Outage of renewable energy units by storms Renewable energy Electricity, heat Preventive, emergency MILP; Gurobi Type-A, Type-D
-[89] Main grid power outages and wind power plant contingencies caused by hurricanes Extreme event starting time Electricity, heat, hydrogen Preventive, emergency Rolling-horizon MILP; Gurobi Type-A, Type-D
-[72] Line outages caused by hurricane Failed transmission lines, their failure, and repair times Electricity, hydrogen Preventive, emergency Stochastic multi-objective MILP Type-D
-[23] Line outages and reduced PV generation caused by freezing rain diaster Line outages Electricity, heating, cooling, hydrogen Preventive, emergency MILP; Gurobi Type-A, Type-D
-[11] Line outages caused by extreme weather Line faults Electricity Preventive, emergency Two stage robust optimization; C&CG Type-B, Type-F
-[73] Line outages caused by hurricane Damaged transmission lines, hurricane time, and repair time Electricity, heat, hydrogen Preventive, emergency Stochastic MILP Type-C, Type-D
-[77] Main grid disconnection and line outages caused by extreme events Line faults Electricity Preventive, emergency Tri-level MILP; Nested C&CG Type-C, Type-F
-[81] Main grid power outages caused by extreme events Load, renewable energy, and travelling time Electricity, heat Preventive, restoration Stochastic MILP; Gurobi Type-D
-[87] Lines/equipment failures incurred by hurricanes Lines/component failure probability Electricity Preventive, restoration Stochastic multi-objective programming; NSGA-II Type-C, Type-J
-[75] Line outages caused by typhoons $\times$ Electricity, hydrogen Preventive, restoration MISOCP; CPLEX Type-A, Type-J, Type-K
-[122] Main grid disconnection during emergencies $\times$ Electricity, heat, cooling, hydrogen Emergency, restoration MILP; CPLEX Type-G, Type-J
-[21] Line outage caused by extreme events Renewable generation, load demands, and charging level of EVs Electricity, heat, cooling, hydrogen Emergency, restoration Bi-level MILP; Adaptive ADMM algorithm Type-G, Type-J
-[123] Line outages and insufficient gas supply caused by extreme events Demand and renewable generation Electricity, heat, gas, hydrogen Emergency, restoration Stochastic MIQCP; Augmented ADMM algorithm Type-F, Type-D, Type-J
-TABLE XI: Representative Hydrogen-related Operation Measures for HMES resilience enhancement
+TABLE XII: Summary of Operation Enhancement for HMES (Multi-stage)
+Ref. Parameter/event uncertainty Contingency uncertainty Uncertainty processing methods Operation stages Problem types and solving methods Hydrogen-related measure types
+[87] $\times$ NkN-k line contingencies Scenarios Preventive, emergency Stochastic MISOCP; Receding horizon method A, D
+[83] Renewable energy Outage of renewable energy units Box uncertainty set Preventive, emergency MILP; Gurobi A, D
+[90] Extreme event starting time Deterministic wind generation curtailment Scenarios Preventive, emergency Rolling-horizon MILP; Gurobi A, D
+[72] $\times$ Failed transmission lines, their failure, and repair times Scenarios Preventive, emergency Stochastic multi-objective MILP D
+[23] $\times$ Component outages Scenarios Preventive, emergency MILP; Gurobi A, D
+[11] $\times$ Line faults Uncertainty sets Preventive, emergency Two stage robust optimization; C&CG B, F
+[73] $\times$ Line outages Scenarios Preventive, emergency Stochastic MILP C, D
+[77] $\times$ NkN-k line contingencies Uncertainty sets Preventive, emergency Tri-level MILP; Nested C&CG C, F
+[82] Load, renewable energy, and travelling time Deterministic grid outages Scenarios Preventive, restoration Stochastic MILP; Gurobi D
+[88] $\times$ Lines/component failure Scenarios Preventive, restoration Stochastic multi-objective programming; NSGA-II C, J
+[75] $\times$ Deterministic line outage $\times$ Preventive, restoration MISOCP; CPLEX A, J, K
+[138] $\times$ Deterministic grid outage $\times$ Emergency, restoration MILP; CPLEX G, J
+[21] Renewable generation, load demands, EV charging level Deterministic line outage Scenarios Emergency, restoration Bi-level MILP; Adaptive ADMM algorithm G, J
+[139] $\times$ Deterministic line outages and insufficient gas supply $\times$ Emergency, restoration Stochastic MIQCP; Augmented ADMM algorithm F, D, J
+TABLE XIII: Representative Hydrogen-related Operation Measures for HMES resilience enhancement
 Stages Measure types Hydrogen-related measures Examples of specific measures
-Preventive Type-A HESS pre-charging Increasing HESS level for reserve[50][75][82][84, 85, 86][23]
-Type-B HTT scheduling HRS-based microgrid formation assisted by transport-based measures (e.g., HTTs and mobile hydrogen trucks) [30], pre-disaster HTT scheduling[11]
+Type-A HESS pre-charging Increasing HESS level for reserve[50][75][83][85, 86, 87][23]
+Preventive Type-B HTT scheduling HRS-based microgrid formation assisted by transport-based measures (e.g., HTTs and mobile hydrogen trucks) [30], pre-disaster HTT scheduling[11]
 Type-C MHER location allocation Location allocation for HTTs[77], sizing and locating for HFCVs[88]
-Emergency Type-D HESS discharging Hydrogen tank-driven fuel cell for independent power supply during outages[51], discharging batteries and HESS[89], HESS and CHP work at maximum power[82], discharging HESS for serving HFCVs[23], discharging HESS for power supply[86]
-Type-E Natural gas pipeline-assisted hydrogen use, hydrogen pipeline scheduling Utilizing hydrogen enriched natural gas for fuel support[31], using pipelines as large-scale energy storage[119], using hydrogen stored in the pipeline for power support[117]
-Type-F HTT/hydrogen truck scheduling Using mobile hydrogen trucks for critical load support[18][19], fuel cell power supply supported by hydrogen from HTTs[11], dynamic re-routing of HTTs[77]
+Type-D HESS discharging Hydrogen tank-driven fuel cell for independent power supply during outages[51], discharging batteries and HESS[90], HESS and CHP work at maximum power[83], discharging HESS for serving HFCVs[23], discharging HESS for power supply[87]
+Type-E Natural gas/hydrogen pipeline-assisted energy supply Utilizing hydrogen enriched natural gas for fuel support[31], using pipelines as large-scale energy storage[136], using hydrogen pipeline for power support[134]
+Emergency Type-F HTT/hydrogen truck scheduling Using mobile hydrogen trucks for critical load support[18][19], fuel cell power supply supported by hydrogen from HTTs[11], dynamic re-routing of HTTs[77]
 Type-G Hydrogen demand adjustment Shedding non-critical hydrogen demand[136], restoring the integrity of compromised hydrogen demand data [59], HRSs adjust the charging schedule of HFCVs[21][138]
 Type-H Hydrogen transformation Using hydrogen generated from renewables for synthetic natural gas production to satisfy demand[76]
-Restoration Type-I Mobile hydrogen generation and storage Combined operation of mobile wind turbines and HESSs[80]
-Type-J HFCV or hydrogen truck scheduling HFCV routing and dispatch[75][76][87], HFCV for power support[75], hydrogen fuel cell trucks are scheduled to HRSs for charging and dispatched to connection points for power supply[21][122]
-Type-K HESS scheduling Scheduling HESS with other components[81], HESS discharging[74][75]
+Type-I Mobile hydrogen storage Combined operation of mobile wind turbines and HESSs[81]
+Type-J HFCV scheduling HFCV routing and dispatch[75][78][88], HFCV for power support[75]
+Restoration Type-K HESS scheduling Scheduling HESS with other components[82], HESS discharging[74][75]
 Type-L Hydrogen fuel cell scheduling Hydrogen fuel cells as black-start resources[57]
@@
 Simulation results indicated that the proposed approach can harness maximum capacities from smart prosumers for reducing load shedding and improving the infrastructure resilience.
-Figure 15: The framework of resilient scheduling for electricity-hydrogen distribution networks
-Summary: For easy understanding of existing works related to resilient operation for HMESs, we summarize their characteristics in Table IX and Table X, and list hydrogen-related measures in each response stage in Table XI.
+Figure 15: The framework of resilient scheduling for electricity-hydrogen distribution networks[77]
+Summary: For easy understanding of existing works related to resilient operation for HMESs, we summarize their characteristics in Table XI and Table XII, and list hydrogen-related measures in each response stage in Table XIII.
 It can be observed that hydrogen plays an important role in every response stage, such as pre-charging HESS in the preventive response stage, hydrogen for fuel support via natural gas pipeline in the emergency response stage, and scheduling MHERs in the restoration response stage.
@@
 By exploiting such elasticity, many resilience enhancement measures related to hydrogen pipelines could be adopted.
-For example, MESs can temporarily store hydrogen in the pipeline when there are extra renewable energy but limited hydrogen tank space.
-When extreme event happens, the hydrogen pipeline can release hydrogen to drive fuel cell for energy supply.
-Such linepack utilization is helpful for enhancing resilience of MESs[125].
+For example, MESs can temporarily store hydrogen in the pipeline when there is surplus renewable energy but limited hydrogen tank space.
+When an extreme event happens, the hydrogen pipeline can release hydrogen to drive fuel cells for energy supply.
+Such linepack utilization is helpful in enhancing resilience of MESs[141].
 Moreover, during extreme events (such as main grid outages or natural gas supply interruptions), hydrogen pipelines can be directionally dispatched to recover hydrogen supply to critical loads such as hospitals and communication base stations.
@@
 A possible solution is to develop comprehensive resilience metrics by considering hydrogen-specific attributes (e.g., hydrogen system availability ratio and pressure recovery time), load loss, the number of customers affected, and carbon resilience jointly.
-VIII-B Temporally-correlated Event-targeted Scenario Generation for Planning
-In existing works on planning methods for the resilience enhancement of HMESs, uncertainties related to system parameters (e.g., distributed generation output and versatile demands) are seldom considered, or just several representative days are considered.
-In fact, the contingency duration incurred by an extreme event (e.g., extreme snow and freezing ice disasters) may last for several days[10][131].
-Thus, the representative day is not long enough to represent the contingency uncertainty.
-Moreover, when taking system parameters into consideration, their correlations should be utilized to generate more representative and practical scenarios.
-For example, in some places, when heat demand is needed in winter, the corresponding cold demand may be not required.
-Similarly, when cold demand is needed in summer, the corresponding heat demand may be negligible[132].
-In addition, extreme events generally have spatial features and their uncertainties should also be considered for scenario generation[47].
-In summary, how to generate temporally-correlated and event-targeted scenarios for planning method design requires further investigation.
-A possible solution for generating temporally-correlated scenarios is to generate normalized multi-parameter vector scenarios according to the duration of extreme events and condense them using scenario reduction methods.
-Finally, the condensed normalized multi-parameter vectors are recovered using denormalization operations.
+VIII-B Extreme-event Scenario Generation for HMES Resilience Planning
+In existing studies on HMES resilience planning, only a limited number of works have explicitly considered extreme-event uncertainty.
+Among them, extreme-event scenario generation has mainly been conducted using generative adversarial networks (GANs) [47] and climate- and weather-model-driven methods [105].
+GAN-based approaches are attractive because they can learn high-dimensional spatio-temporal patterns from data and generate large numbers of synthetic extreme scenarios, which is particularly useful for HMESs where prolonged renewable deficits may coincide with heating or cooling stress.
+Climate- and weather-model-driven methods, by contrast, can provide physically grounded realizations of persistent meteorological extremes, such as prolonged heatwaves, droughts, and renewable-energy droughts, which are often insufficiently represented in historical observations.
+Despite these advantages, both approaches have important limitations, as summarized in Table X.
+In particular, GAN-based methods are typically data-hungry and may generate physically implausible scenarios in the absence of sufficient domain constraints.
+Climate- and weather-model-driven methods, on the other hand, suffer from high computational burden, the need for downscaling and bias correction, the difficulty of translating raw climate outputs into planning-ready scenarios for HMESs, and structural uncertainty arising from imperfect or simplified model formulations.
+More importantly, these limitations are not confined to the above two categories, but are also relevant to other extreme-event scenario generation methods listed in Table X.
+Among them, one of the most critical unresolved issues is extreme-data scarcity.
+By definition, severe extreme events are rare, while compound extremes are even less frequently observed.
+As a result, the available data record is often too short to support robust characterization of long-duration, spatially correlated, and concurrent extreme events.
+This challenge is particularly pronounced for HMESs, because the resilience value of hydrogen is expected to emerge precisely under low-frequency but high-impact events, whereas such conditions are poorly represented in ordinary operational datasets.
+Therefore, addressing extreme-data scarcity is a key prerequisite for more credible resilience-oriented planning.
+Several promising directions may help address this challenge.
+First, extreme-value-theory-assisted deep generative models offer a valuable hybrid framework, in which deep generative models capture high-dimensional dependence structures while extreme value theory improves the representation and extrapolation of tail behavior [125].
+Second, physics-informed generative models provide a way to enhance the physical plausibility of synthetic extreme scenarios by embedding conservation laws, known meteorological drivers, or other domain constraints into the generation process [125].
+Third, response-aware scenario generation methods can identify representative climate scenarios according to system response, thereby reducing scenario complexity while preserving the most critical stress patterns for resilience-oriented planning [132].
+These directions suggest that future extreme-event scenario generation for HMES resilience planning should move beyond purely data-fitting approaches toward hybrid frameworks that are tail-aware, physically consistent, and decision-relevant.
 VIII-C Multi-type Temporal-spatial Cyber-physical Contingency Modeling under Compound Extreme Events
@@
 Such delay may span from minutes to hours.
-In addition to slow dynamics, hydrogen pipeline can be used to store hydrogen temporarily due to the compressibility of hydrogen, which can help to absorb fluctuations in renewable energy generation or load demand without the immediate need for additional storage tanks.
-Compared with hydrogen tank/pipelines, large-scale hydrogen energy storage systems (e.g., salt-cavern hydrogen storage) have slower dynamics, which are used as SHSs to adjust the weekly/monthly source-load imbalance[54].
-In contrast, to capture the minute-level fluctuations of renewable energy and load demand, electricity system could be scheduled at minute timescale.
+In addition to slow dynamics, hydrogen pipelines can be used to store hydrogen temporarily due to the compressibility of hydrogen, which can help to absorb fluctuations in renewable energy generation or load demand without the immediate need for additional storage tanks.
+Compared with hydrogen tanks/pipelines, large-scale hydrogen energy storage systems (e.g., salt-cavern hydrogen storage) have slower dynamics, which are used as SHSs to adjust the weekly/monthly source-load imbalance[54].
+In contrast, to capture the minute-level fluctuations of renewable energy and load demand, electricity system could be scheduled on a minute timescale.
 Thus, multi-timescale modeling should be considered in the planning and operation of hydrogen-enabled MESs.
@@
 is attributed to heat demand.
-When HMESs with versatile energy demands are considered, more strong couplings of power/gas/heat networks would be incurred.
+When HMESs with versatile energy demands are considered, more strong couplings of power/gas/heat networks would arise.
 To manage such systems of systems efficiently, multi-network coordination is required.
@@
 VIII-F LLM-assisted Whole-Process Resilience Enhancement for HMESs
-LLMs (e.g., GPT-4) have demonstrated remarkable capabilities in understanding, synthesizing, and reasoning across vast domains of unstructured and structured information[149, 150].
+LLMs (e.g., GPT-series models) have demonstrated remarkable capabilities in understanding, synthesizing, and reasoning across vast domains of unstructured and structured information[164, 165].
 Their contextual awareness, semantic reasoning, and adaptability position them as promising tools for supporting the whole-process resilience enhancement of HMESs.
 Firstly, the current resilience metrics failed to capture the hydrogen-related attributes, e.g., hydrogen availability, tank/pipeline pressure recovery, and sector-coupled response.
-In contrast, LLMs can extract, analyze, and generate multi-dimensional, cross-sectoral resilience indicators by mining technical literature, policy documents, and incident reports, providing a foundation for standardizing hydrogen-specific evaluation frameworks.
-Secondly, contingency modeling under extreme events in HMESs is hindered by the complexity of hydrogen-related subsystems, while LLMs can contribute to contingency modeling by constructing structured fault trees and scenario libraries based on natural language descriptions, enabling the automatic generation of probabilistic or rule-based failure scenarios under diverse stress conditions.
-Thirdly, from a planning enhancement perspective, LLMs can support scenario generation and strategic siting of hydrogen assets by reasoning over multi-objective (e.g., operation cost, carbon emission, and resilience) trade-offs.
-Finally, from the operation enhancement perspective, LLMs offer potential in assisting the dynamic dispatch and routing of MHERs since they can interpret complex, multi-modal inputs (ranging from sensor data to emergency protocols) and contribute to human-in-loop decision.
+In contrast, LLMs can help extract and organize hydrogen-specific resilience attributes from technical literature, safety standards, policy documents, and incident reports, thereby supporting the development of more comprehensive resilience metrics for HMESs.
+Secondly, contingency modeling under extreme events in HMESs is hindered by the complexity of hydrogen-related subsystems.
+For example, during a typhoon event, transmission outages may coincide with road blockages, temporary natural gas shortages, and constrained hydrogen delivery, creating cascading impacts that are difficult to model using only predefined structured data.
+In this context, LLMs can assist contingency modeling by converting natural-language descriptions from operation manuals, fault logs, and emergency response documents into structured fault trees, event chains, and scenario libraries, which can improve the efficiency of constructing probabilistic or rule-based contingency models under diverse stress conditions.
+Thirdly, from a planning perspective, LLMs can support semantically grounded extreme-event scenario generation since their training are conducted based on massive textual corpora[125].
+In addition, LLMs can support strategic siting of hydrogen-related assets by transforming multi-source textual knowledge into candidate rules, constraints, and evaluation criteria.
+For instance, an LLM may infer that a location close to wind power resources is economically attractive for electrolysis, but unsuitable for emergency hydrogen supply if it is far from hospitals, communication hubs, or major transport corridors.
+Finally, from an operational perspective, LLMs can support the dynamic dispatch of HMESs by integrating heterogeneous real-time information (e.g., sensor readings, traffic updates, weather alerts, and emergency protocols) and assisting decision-making on hydrogen allocation, mobile resource routing, and recovery prioritization.
+Therefore, rather than replacing physics-based models or optimization tools, LLMs can serve as an intelligent interface that connects fragmented knowledge, planning rules, and real-time operational information, thereby enhancing the resilience of HMESs across assessment, modeling, planning, and operation.
 IX Conclusions
@@
 Under the proposed framework, the widely used resilience metrics and event-oriented contingency models are summarized.
-Then, it classifies the hydrogen-enabled planning measures for resilience enhancement of HMESs according to the type of hydrogen-related facilities and summarizes the typical planning problem formulation frameworks.
+Then, it classifies the hydrogen-enabled planning measures for resilience enhancement of HMESs according to the type of hydrogen-related facilities and summarizes the classifications of uncertainty processing methods and scenario generation methods, as well as planning problem formulation frameworks.
 Next, it categorizes the hydrogen-enabled operation measures for the resilience enhancement of HMESs according to three types of operation response stages and summarizes the role of hydrogen in each response stage.
-Finally, several research gaps and future directions are identified, i.e., (1) Comprehensive HMES resilience metrics with the consideration of hydrogen-specific attributes (e.g., hydrogen system availability ratio and pressure recovery time), load loss, the number of customers affected, and carbon resilience should be designed; (2) Generating temporally-correlated event-targeted scenario generation for planning based on normalized multi-parameter vector scenarios and scenario reduction methods; (3) Achieving multi-type temporal-spatial contingency modeling under compound extreme events based on spatially-aware dynamic Bayesian networks and Petri-net; (4) Implementing multi-network multi-timescale coordination planning and operation based on hierarchical MPC or hierarchical deep reinforcement learning; (5) Designing advanced artificial intelligence-based multi-objective optimization methods for the planning and operation of low-carbon and resilient HMES; (6) Adopting LLM-assisted whole-process resilience enhancement for HMESs in aspects of proper resilience metrics, event-oriented contingency modeling, scenario generation and smart decision in planning process, and MHER scheduling and routing in operation process.
+Finally, several research gaps and future directions are identified, i.e., (1) Comprehensive HMES resilience metrics with the consideration of hydrogen-specific attributes (e.g., hydrogen system availability ratio and pressure recovery time), load loss, the number of customers affected, and carbon resilience should be designed; (2) Generating tail-extrapolation-aware, physical-informed, response-relevant extreme-event scenarios for HMES resilient planning based on extreme value theory-assisted deep generative models, physics-informed generative models, and response-aware scenario generation methods; (3) Achieving multi-type temporal-spatial contingency modeling under compound extreme events based on spatially-aware dynamic Bayesian networks and Petri-net; (4) Implementing multi-network multi-timescale coordination planning and operation based on hierarchical MPC or hierarchical deep reinforcement learning; (5) Designing advanced artificial intelligence-based multi-objective optimization methods for the planning and operation of low-carbon and resilient HMES; (6) Adopting LLM-assisted whole-process resilience enhancement for HMESs in aspects of proper resilience metrics, event-oriented contingency modeling, scenario generation and smart decision in planning process, and MHER scheduling and routing in operation process.
 While the above-mentioned technical strategies are essential for enhancing the resilience of HMESs, business and market models also play a crucial role.